### v1

# SoundMorpher: Perceptually-Uniform Sound Morphing with Diffusion Model

Abstract
We present SoundMorpher, a sound morphing method that generates perceptually uniform morphing trajectories using a diffusion model. Traditional sound morphing methods models the intractable relationship between morph factor and perception of the stimuli for resulting sounds under a linear assumption, which oversimplifies the complex nature of sound perception and limits their morph quality. In contrast, SoundMorpher explores an explicit proportional mapping between the morph factor and the perceptual stimuli of morphed sounds based on Mel-spectrogram. This approach enables smoother transitions between intermediate sounds and ensures perceptually consistent transformations, which can be easily extended to diverse sound morphing tasks. Furthermore, we present a set of quantitative metrics to comprehensively assess sound morphing systems based on three objective criteria, namely, correspondence, perceptual intermediateness, and smoothness. We provide extensive experiments to demonstrate the effectiveness and versatility of SoundMorpher in real-world scenarios, highlighting its potential impact on various applications such as creative music composition, film post-production and interactive audio technologies. See SoundMorpher-demo for the listening demonstration.
1 Introduction
Sound morphing is a technique to create a seamless transformation between multiple sound recordings. The goal is to produce perceptual intermediate sounds that gradually change from one sound to another. Sound morphing has a wide range of applications, including music compositions, synthesizers, psychoacoustic experiments to study timbre spaces (Caetano & Rodet, 2011; Hyrkas, 2021), and practical applications such as film post-production, AR or VR interactive games, and adaptive audio content in video games (Qamar et al., 2020; Siddiq, 2015).
Traditional sound morphing methods used the interpolation principle of sound synthesis technique, which relies on interpolating the parameters of a sinusoidal model (Tellman et al., 1995; Osaka, 1995; Williams et al., 2014). Others make use of digital signal processing techniques to explore high-level audio features in the time-frequency domain to achieve more effective and continuous morphing (Williams et al., 2014; Brookes & Williams, 2010; Caetano & Rodet, 2010; 2011; Roma et al., 2020; Caetano, 2019). However, these methods are limited to applications such as producing inharmonic and noisy environmental sounds (Gupta et al., 2023; Kamath et al., 2024). Despite the increasing interest in applying machine learning to sound generation, there has only been limited exploration in sound morphing. Recent approaches (Zou et al., 2021; Gupta et al., 2023; Kim et al., 2019b; Kamath et al., 2024) have shown their superior effectiveness compared to traditional methods in various scenarios. However, we observed several critical limitations of those existing methods. Firstly, they are primarily designed for static or cyclostationary morphing (see Sec. 3.1), limiting their applicability to dynamic sound transformations. Secondly, these approaches often lack sufficient quantitative evaluation, limiting further analysis of their effectiveness. Thirdly, they require training on task-specific datasets, which limits their application in different scenarios. Most importantly, they typically assume a linear relationship between morphing factors and sound perception, and achieve smooth morphing by gradually changing the morph factor. This assumption oversimplifies the complex nature of sound perception, as gradually changing morph factors does not inherently result in smooth perceptual transitions. To this end, our goal is to develop a method that achieves perceptually coherent morphing, ensuring seamless and natural sound transition.
In this paper, we introduce SoundMorpher, a sound morphing method that produces perceptually smooth and intermediate morphing, comprising the following key contributions.
  • SoundMopher is the first open-world sound morphing method based on a pre-trained diffusion model, which integrates typical morph tasks such as static, dynamic and cyclostationary morphing. Unlike prior works (Kim et al., 2019b; Gupta et al., 2023), SoundMorpher can be broadly applied to various real-world tasks without requiring extensive retraining.
  • We propose the sound perceptual distance proportion (SPDP), which explicitly connects morph factors and perceptual stimuli of morphed results. This allows SoundMorpher to produce morphing paths with a uniform change in perceptual stimuli, achieving more seamless perceptual transitions compared to existing methods (Kamath et al., 2024).
  • We adapt a set of comprehensive quantitative metrics according to criteria proposed by Caetano & Osaka (2012) for evaluation, addressing the lack of quantitative assessment for sound morphing systems (Caetano, 2019; Zou et al., 2021; Caetano & Rodet, 2013) and may offer insights for analyzing and comparing future sound morphing methods.
  • We provide extensive experiments to demonstrate that SoundMorpher can be effectively applied to several potential applications in broader real-world scenarios, including musical instrument timbre morphing, music morphing and environmental sound morphing.
2 Related Work
In this section, we first present a detailed review of related works on sound morphing task. Then, we also briefly introduce tasks that are similar to sound morphing and clarify the differences.
Sound morphing. Traditional sound morphing methods rely on interpolating parameters of a sinusoidal sound synthesis model (Tellman et al., 1995; Osaka, 1995; Williams et al., 2014; Primavera et al., 2012). To achieve more effective and continuous morphing, Williams et al. (2014); Brookes & Williams (2010); Caetano & Rodet (2010; 2011); Roma et al. (2020); Caetano (2011) target on exploring perceptual spectral domain audio features by digital signal processing techniques, such as MFCCs, spectral envelope, etc.. Others such as Kazazis et al. (2016) involve a hybrid approach that extracts audio descriptors to morph accordingly and interpolate between the spectrotemporal fine structures of two endpoints according to morph factors. Machine learning sound morphing methods offer advantages such as high morphing quality by leveraging semantic representation interpolation within a model instead of traditional audio feature interpolation. Zou et al. (2021) proposes a non-parallel many-to-one static timbre morphing framework that integrates and fine-tunes the machine learning technique (i.e., DDSP-autoencoder (Engel et al., 2020)) with spectral feature interpolation (Caetano & Rodet, 2013). Kim et al. (2019b) targets synthesizing music corresponding to a note sequence and timbre, which uses non-linear instrument embedding as timbre control parameters under a pretrained WaveNet (Engel et al., 2017) to achieve timbre morphing between instruments. Luo et al. (2019) learns latent distributions of VAEs to disentangle representations for pitch and timbre of musical instrument sounds. Tan et al. (2020) uses a GM-VAE to achieve style morphing to generate realistic piano performances in the audio domain following temporal style conditions for piano performances, which morphs the conditions such as onset roll and MIDI note into input audio. MorphGAN (Gupta et al., 2023) targets on audio texture morphing by interpolating within conditional parameters, and trained the model on a water-wind texture dataset. A recent concurrent work by Kamath et al. (2024) uses a pre-trained AudioLDM (Liu et al., 2023) to morph sound between two text prompts. In contrast, we focus on classical sound morphing, where the morphing process is performed directly between two given audios rather than between text prompts. A key advantage of our method is its ability to provide precise guidance during the morphing process, as the target audio delivers exact information on how the source sound should evolve—something that text prompts cannot always achieve, for example, morphing between two music compositions.
Synthesizer preset interpolation. Synthesizer preset interpolation achieves sound morphing by developing models that compute interpolations within the domain of synthesis parameters for a black-box synthesizer (Le Vaillant & Dutoit, 2023; Dutoit et al., 2023; Le Vaillant & Dutoit, 2024). Unlike classical sound morphing, which perceptually blends two audio files into an intermediate sound, synthesizer preset interpolation treats the synthesizer as a non-differentiable black box, with presets composed of both numerical and categorical parameters. By smoothly interpolating between these presets, the task aims to achieve seamless morphing of synthesized sounds.
Text-to-audio editing. Text-to-audio editing is the process of using text queries to edit audio. With the success of diffusion models in image editing tasks, recent works target zero-shot audio editing with text instructions (Manor & Michaeli, 2024; Zhang et al., 2024; Lan et al., 2024) involving tasks such as inpainting, outpainting, timbre transfer, music genre transfer, or vocals removal.
Timbre transfer. Timbre transfer is a specific task that aims at converting the sound of a musical piece by one instrument (source) into the same piece played by another instrument (target). This concerns the task of converting a musical piece from one timbre to another while preserving the other music-related characteristics (Comanducci et al., 2024; Jain et al., 2020; Li et al., 2024).
Voice conversion and morphing. Voice conversion (VC) involves modifying vocal characteristics of a source speech to match a target speaker, either by using target speeches or text (Li et al., 2023; Yao et al., 2024; Niu et al., 2024; Sheng et al., 2024). The primary objective of VC is to alter the vocal identity to closely resemble the target voice style, while preserving the linguistic content of the source speech. Voice morphing is a broader scope, focusing on blending or transforming one voice into another. This often involves creating an intermediate voice that incorporates characteristics of both source and target voices, allowing for gradual transitions between them (Sheng et al., 2024).
3 Preliminaries
3.1 Sound Morphing
Sound morphing aims to produce intermediate sounds as different combinations of model source sound $\hat{S}_{1}$ and target sound $\hat{S}_{2}$ (Caetano & Rodet, 2010; 2011), which can be formulated as
$M(\alpha,t)=(1-\alpha(t))\hat{S}_{1}+\alpha(t)\hat{S}_{2}$ (1)
Each step is characterized by one value of a single parameter $\alpha$, the so-called morph factor, which ranges between 00 and $1$, where $\alpha=0$ and $\alpha=1$ produce resynthesized source and target sounds, respectively. Due to the intrinsic temporal nature of sounds, sound morphing usually involve three main types: dynamic morphing, where $\alpha$ gradually transfers from 00 to $1$ over time dimension (Kazazis et al., 2016), static morphing, where a single morph factor $\alpha$ leads to an intermediate sound between source and target (Sethares & Bucklew, 2015), and cyclostationary morphing where several hybrid sounds are produced in different intermediate points (Slaney et al., 1996).
To solve the limitation on previous works that target on expensive perceptual evaluation only, Caetano & Osaka (2012) proposes three objective criteria for sound morphing techniques: (1) Correspondence. The morph is achieved by a description whose elements are intermediate between source and target sounds, highlighting semantic level transition; (2) Intermediateness. The morphed objects should be perceived as intermediate between source and target sounds, evaluating perceptual level correlation; (3) Smoothness. The morphed sounds should change gradually (i.e., ‘smoothly’) from source to target sounds, by the same amount of perception increment. Under the assumption of linear perceptual stimuli, adding the same factor should increase the same amount of perception. In this study, we evaluate SoundMorpher according to the three criteria by a series of comprehensive objective quantitative metrics in addition to perceptual evaluation.
3.2 Latent Diffusion Model on Audio Generation
SoundMorpher utilizes a pretrained text-to-audio (TTA) latent diffusion model (LDM) (Rombach et al., 2022) to achieve sound morphing. This approach offers the advantage of performing various types of sound morphing without the need to train the entire model or use additional datasets. Specifically, we use AudioLDM2 (Liu et al., 2024), a multi-modality conditions to audio model. It employs a pre-trained variational autoencoder (VAE) (Kingma & Welling, 2013) to compress audio $x$ into a low-dimension latent space as VAE representations $z$. AudioLDM2 generates latent variables $z_{0}$ from a Gaussian noise $z_{T}$ given the condition $C$ and further reconstruct audio $\hat{x}$ from $z_{0}$ by VAE decoder and a vocoder (Kong et al., 2020). AudioLDM2 uses an intermediate feature $Y$ as an abstraction of audio data $x$ to bridge the gap between conditions $C$ and audio $x$, named language of audio (LOA). The LOA feature is obtained by a AudioMAE (Huang et al., 2022; Tan et al., 2024) and a series of post-processing formulated as $Y=\mathcal{A}(x)$. The generation function $\mathcal{G}(\cdot)$ is achieved by a LDM. In the inference phase, AudioLDM2 approximates LOA feature by the given condition as $\hat{Y}=\mathcal{M}(C)$ using a fine-tuned GPT-2 model (Radford et al., 2019). Then generates audios conditioned on the estimated LOA feature $\hat{Y}$ and an extra text embedding $E_{T5}$ from a FLAN-T5 (Chung et al., 2024) with a LDM as $\hat{x}=\mathcal{G}(\hat{Y},E_{T5})$. We denote the conditional embeddings in AudioLDM2 as $E=\{\hat{Y},E_{T5}\}$, therefore, the generative process becomes $\hat{x}=\mathcal{G}(E)$.
Diffusion Models. The LDM performs a forward diffusion process during training, which is defined as a Markov chain that gradually adds noise to the VAE representation $z_{0}$ over $T$ steps as $z_{t}=\sqrt{1-\beta_{t}}z_{t-1}+\sqrt{\beta_{t}}\epsilon_{t}$. where $\epsilon_{t}\sim N(0,I)$ and noise schedule hyperparameter $\beta_{t}\in[0,1]$. Therefore, we can derive the distribution of $z_{t}$ given $z_{0}$ as $q(z_{t}|z_{0})=\sqrt{\gamma_{t}}z_{0}+\sqrt{1-\gamma_{t}}\epsilon_{t}$, where $\gamma_{t}=\prod_{t=1}^{t}1-\beta_{t}$. The LDM learns a backward transition $\epsilon_{\theta}(z_{t},t)$ from the prior distribution $N(0,I)$ to the data distribution $z$, that predicts the added noise $\epsilon_{t}$ (Ho et al., 2020). Following the objective function of denoising diffusion probabilistic models (DDPM) (Ho et al., 2020), the objective function for training AudioLDM2 is
$\text{min}_{\theta}\mathcal{L}_{DPM}=\text{argmin}_{\theta}[\mathbb{E}_{z_{0},E,t\sim\{1,...,T\}}||\epsilon_{\theta}(z_{t},E,t)-\epsilon_{t}||_{2}^{2}]$ (2)
To reduce computational demands on inference, AudioLDM2 uses denoising diffusion explicit models (DDIM) (Song et al., 2020), which provides an alternative solution and enables significantly reduced sampling steps with high generation quality. The DDIM reverse diffusion process is
$z_{t-1}=\sqrt{\gamma_{t-1}}(\frac{z_{t}-\sqrt{1-\gamma_{t}}\epsilon_{\theta}(z_{t},E,t))}{\sqrt{\gamma_{t}}})+\sqrt{1-\gamma_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(z_{t},E,t)+\sigma_{t}\epsilon_{t}$ (3)
We can revise a deterministic mapping between $z_{0}$ and its latent state $z_{T}$ once the model is trained (Dhariwal & Nichol, 2021; Yang et al., 2023) by the following equation
${}\frac{z_{t+1}}{\sqrt{\gamma_{t+1}}}-\frac{z_{t}}{\sqrt{\gamma_{t}}}=(\sqrt{\frac{1-\gamma_{t+1}}{\gamma_{t+1}}}-\sqrt{\frac{1-\gamma_{t}}{\gamma_{t}}})\epsilon_{\theta}(z_{t},E,t)$ (4)
4 Method
Given a source and target audio pair $\{x^{(0)},x^{(1)}\}$, sound morphing aims to generate intermediate sounds $x^{(\alpha(t))}$ between the audio pair given morph factors $\alpha\in[0,1]$. To account for the variation of $\alpha$ over time in Equation 1, we discretize the function $\alpha(t)$ where $t\in[0,T]$ into $N$ elements, resulting in a morphed sequence of sounds $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ based on $\{\alpha_{i}\}^{N}_{i=1}$. According to the smoothness criteria proposed by Caetano & Osaka (2012), the desired sound morphing technique should have smooth linear perceptual stimuli when the morph factor $\alpha$ varies in the sequence $\{\alpha_{i}\}^{N}_{i=1}$. Therefore, we define $p_{i}$ to represent the perceptual stimuli of the morphed audio $x^{(\alpha_{i})}$ given morph factor $\alpha_{i}$. However, the relationship $\mathcal{P(\cdot)}$ between morph factor $\alpha$ and perceptual stimuli $p$ is intractable. Our goal is to find a discrete morph factor sequence $\{\alpha_{i}\}^{N}_{i=1}$ such that for each transition, the perceptual stimuli difference $\Delta p$ is a constant value. Therefore, we formulate the problem as
$p_{i+1}-p_{i}\equiv\mathcal{P}(x^{(\alpha_{i+1})})-\mathcal{P}(x^{(\alpha_{i})})=\Delta p,~{}~{}i\in[1,...,N-1]~{}$ (5)
This formulation is a refined sound morphing problem where, rather than controlling morph factor $\alpha$, we control the constant perceptual stimuli difference $\Delta p$ to find the optimal trajectory with morph factors $\{\alpha_{i}\}^{N}_{i=1}$ that will achieve perceptually uniform sound morphing  [^1] See Appendix 7.1 for overall SoundMorpher pseudo algorithm pipeline and further implementation details..
In Section 4.1 we introduce feature interpolation and model adaption with a pre-trained AudioLDM2. This method allows high-quality intermediate morph results to be obtained by controlling morph factor $\alpha$. To achieve perceptually uniform sound morphing as in Equation 5, we explore an explicit connection $\mathcal{P(\cdot)}$ between perceptual stimuli $p$ and morph factor $\alpha$ in Section 4.2. In Section 4.3, we provide extensions of our method on the different morphing methods discussed in Section 3.1 to show the advantages of perceptually uniform sound morphing.
4.1 Feature interpolation and model adaption
Interpolating optimized conditional embeddings. We first introduce text-guided conditional embedding optimization strategy under a pre-trained AudioLDM2, which retrieves corresponding conditional embeddings $E$ of the given audio data. As mentioned in Section 3.2, AudioLDM2 accepts two conditional inputs: LOA feature $Y$ and text embedding $E_{T5}$. We denote $E=\{Y,E_{T5}\}$ as the overall conditional embedding inputs for AudioLDM2. The LOA feature $Y$ is an abstraction of audio data which is semantically structured, and $E_{T5}$ captures sentence-level of representations. To retrieve corresponding conditional embeddings of the given audio data, we first obtain the latent variables $z_{0}^{(0)}$ and $z_{0}^{(1)}$ of audio $x^{(0)}$ and $x^{(1)}$ from the pre-trained VAE in AudioLDM2 pipeline. We initialize a simple common text prompt (e.g., ‘An audio clip of sound’) as a text guidance condition $C$ to obtain $E$ by GPT-2 encoder and FLAN-T5 encoder in AudioLDM2 pipeline, respectively, as $E^{(0)}$ and $E^{(1)}$. Instead of optimizing the model parameters, we freeze the model parameters and optimize the conditional embedding $E^{(0)}$ and $E^{(1)}$ by the denoising objective function in Equation 2
$E^{(0)}=\text{arg min}_{E}\mathcal{L}_{DPM}(z_{0}^{(0)},E;\theta)\text{ and }E^{(1)}=\text{arg min}_{E}\mathcal{L}_{DPM}(z_{0}^{(1)},E;\theta)$ (6)
The optimized conditional embeddings $E^{(0)}$ and $E^{(1)}$ fully encapsulate the abstract details of audios $x^{(0)}$ and $x^{(1)}$. Due to the semantically structured nature of the conditional embeddings, the conditional distributions $p_{\theta}(z|E^{(0)})$ and $p_{\theta}(z|E^{(1)})$ closely mirror the degree of audio variation between the audio pair. To explore the data distribution that conceptually intermediate between $z^{(0)}$ and $z^{(1)}$, we bridge these two distributions through linear interpolation. Specifically, we define the interpolated conditional distribution as $p_{\theta}(z|E^{(\alpha)}):=p_{\theta}(z|(1-\alpha)E^{(0)}+\alpha E^{(1)})$, where $\alpha\in[0,1]$.
Interpolating latent state. The conditional embedding represents the conceptual abstract of audio data. However, we also wish to smoothly morph the content of the audio pair. Following Song et al. (2020) and Yang et al. (2023), we smoothly interpolate between $z_{0}^{(0)}$ and $z_{0}^{(1)}$ by spherical linear interpolation (slerp) to their starting noise $z_{T}^{(0)}$ and $z_{T}^{(1)}$ and further obtained the interpolated latent state $z_{T}^{(\alpha)}:=\frac{\text{sin}(1-\alpha)\omega}{\text{sin}\omega}z_{T}^{(0)}+\frac{\text{sin}\alpha\omega}{\text{sin}\omega}z_{T}^{(1)}$, where $\omega=\text{arccos}(\frac{z_{T}^{(0)\intercal}z_{T}^{(1)}}{||z_{T}^{(0)}||||z_{T}^{(1)}||})$. The denoised latent variable $z_{0}^{(\alpha)}$ is obtained by applying a diffusion denoising process on the interpolated starting noise $z_{T}^{(\alpha)}$ and conditioning on the interpolated conditional embedding $E^{(\alpha)}$. The final morphed audio result $x^{(\alpha)}$ is obtained from $z_{0}^{(\alpha)}$ by the VAE decoder and a vocoder.
Model adaption. Model adaptation helps to limit the degree of morphed variation by suppressing high-density regions that not related to the given inputs (Yang et al., 2023). We use LoRA (Hu et al., 2021) to inject a small amount of trainable parameters for efficient model adaptation. We fine-tune AudioLDM2 with LoRA trainable parameters using $z^{(0)}$ and $z^{(1)}$. See Appendix 7.2 for details.
4.2 perceptually uniform Sound Morphing
Sound perceptual distance proportion (SPDP). The relationship between morph factor $\alpha$ and perceptual stimuli $p$ is intractable. Our goal is to establish an objective quantitative metric that links $p_{i}$ and $x^{(\alpha_{i})}$ as in Equation 5. This metric should satisfy two key conditions: (1) the output $p$ should increase monotonically as $\alpha$ increases; (2) it should accurately represent perceptual differences between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$, ensuring a smooth transition through intermediate states. Therefore, we propose the sound perceptual distance proportion between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$. We define $p_{i}\in\mathbb{R}^{2}$ as a 2D vector to represent the perceptual proximity of $x^{(\alpha_{i})}$ to both $x^{(0)}$ and $x^{(1)}$. Instead of extracting numerous audio features through traditional signal processing techniques, we use Mel-scaled spectrogram to capture perceptual and semantic information on audio. Mel-spectrogram (Tzanetakis & Cook, 2002) provides a pseudo-3D representation of audio signals, with one axis representing time and the other representing frequency on the Mel scale (Stevens et al., 1937), while the values denote the magnitude of each frequency at specific time points. The advantage of using Mel-spectrogram lies in the Mel filter banks, which map frequencies to equal pitch distances that correspond to how humans perceive sound (Sturm, 2013; Müller, 2015). Denoting $x_{mel}^{(\alpha_{i})}$ as the Mel-spectrogram of audio $x^{(\alpha_{i})}$, the SPDP $p_{i}$ between two endpoint audios $x^{(0)}$ and $x^{(1)}$ given $\alpha_{i}$ is defined as
$p_{i}=[\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}},\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}]$ (7)
Binary search with constant SPDP increment. To produce a perceptually smooth morphing trajectory with a constant perceptual stimuli increment, we use binary search to seek the corresponding $\{\alpha_{i}\}^{N}_{i=1}$ based on a constant $\Delta p$. The target SPDP sequence $\{p_{i}\}_{i=1}^{N}$ is obtained by an interpolation $p_{i}=(1-\frac{i-1}{N-1})p^{(0)}+\frac{i-1}{N-1}p^{(1)}$, where the two endpoints are $p^{(0)}=[0,1]^{T}$ and $p^{(1)}=[1,0]^{T}$. See Algorithm 2 in Appendix 7.3 for detail pseudo algorithm.
4.3 Controllable Sound Morphing with Discrete $\alpha$ Series
By controlling the discrete morph factor sequence $\{\alpha_{i}\}^{N}_{i=1}$ to produce a morphed series $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ with constant $\Delta p$, we achieve three typical morphing methods as follows.
Static morphing. To achieve controllable static morphing, we control the target SPDP point $p$, which represents how the desired output perceptually intermediate between $x^{(0)}$ and $x^{(1)}$. We find the corresponding $\alpha$ value by the binary search with the target $p$ and further obtain a morphed result ${x^{(\alpha)}}$. Pseudocodes for static morphing are in Algorithm 3.
Cyclostationary morphing. To produce $N$ perceptually uniform hybrid sounds between $x^{(0)}$ and $x^{(1)}$, we first obtain $N$ uniform interpolated SPDP points $\{p_{i}\}^{N}_{i=1}$. Then we find corresponding morph factors $\{\alpha_{i}\}_{i=1}^{N}$ and further obtain $N$ morphed results $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ as in Algorithm 4.
Dynamic morphing. Dynamic morphing performs sound morphing over time, but one challenge is that if the morphing path fails to ensure perceptual intermediateness and content correspondence, the resulting sounds may exhibit perceptual discontinuities or unnatural intermediate stages. As in Algorithm 5, we obtain $N$ interpolated target SPDP points $\{p_{i}\}^{N}_{i=1}$ with $\Delta p$. The corresponding morph factors $\{\alpha_{i}\}_{i=1}^{N}$ are determined by binary search with the target SPDP points. Each morphed result $x^{(\alpha_{i})}$ contributes a segment of duration $\frac{T}{N}$, producing an audio segment $\tilde{x}^{(\alpha_{i})}$ according to index $i$. The final audio signal is obtained by concatenating these morphed segments, resulting in
$[x_{0},x_{1},...,x_{T}]=\text{concat}(\tilde{x}^{(0)},\tilde{x}^{(\alpha_{1})},...,\tilde{x}^{(1)})$ (8)
Table 1: Timbral morphing for musical instruments compared to baseline on different instruments.
[…]
Piano $\leftrightarrow$ Guitar SMT 24.73 102.57 1.170 0.116 $\pm$ 0.074 1.263 0.122
Ours 5.21 41.11 0.404 0.044 $\pm$ 0.020 0.466 0.132
Harp $\leftrightarrow$ Kalimaba SMT 13.46 88.89 1.495 0.150 $\pm$ 0.117 1.355 0.182
Ours 4.67 37.92 0.768 0.076 $\pm$ 0.089 0.462 0.159
Taiko $\leftrightarrow$ Hihat SMT 8.51 131.57 2.339 0.234 $\pm$ 0.332 1.584 0.732
Ours 3.32 47.59 1.314 0.131 $\pm$ 0.058 0.359 0.102
Piano $\leftrightarrow$ Violin SMT 21.38 90.63 1.902 0.190 $\pm$ 0.069 0.558 0.217
Ours 3.42 20.14 0.782 0.078 $\pm$ 0.020 0.415 0.085
Piano $\leftrightarrow$ Organ SMT 21.36 63.26 1.291 0.129 $\pm$ 0.074 1.106 0.097
Ours 3.29 19.73 0.233 0.023 $\pm$ 0.010 0.423 0.097
5 Experiment
In this section, we showcase three applications of SoundMorpher in real-world scenarios: Timbral morphing for musical instruments, Environmental sound morphing, and Music morphing.
5.1 Evaluation Metric
[…]
5.2 Timbral morphing for musical instruments
Sound morphing can allow timbral morphing between the sound of two known musical instruments, creating sounds from unknown parts of the timbre space (McAdams, 2013; McAdams & Goodchild, 2017). Timbral morphing for musical instruments involves transitioning between timbres of two different musical instruments to create a new sound. This new sound could possess characteristics of both original timbres as well as new timbral qualities between them, which usually applied to creative arts. In this experiment, we perform timbral morphing for isolated musical instruments given two recordings of the same musical composition played by different music instruments.
Dataset. To ensure high quality of paired composition musical instrument on timbral study, we selected 22 paired musical instrument composition samples from demonstration pages of musical timbre transform projects, MusicMagus (Zhang et al., 2024) and Timbrer (Kemppinen P., 2020). The paired samples have durations varies from 5s to 10s, with 16.0kHz and 44.01kHz, which involve 5 groups of instrument pairs: 2 paired samples of piano-violin; 10 paired samples of piano-guitar; 1 paired sample of taiko-hihat; 1 paired sample of piano-organ, and 8 paired samples of harp-kalimaba.
(a) SMT
(b) SoundMopher (ours)
Figure 1: Timbre space visualization of morph trajectories for piano-organ timbre morphing. Compared to SMT, SoundMorpher produces a smoother and continuous morph in the timbre space.
Baseline. We compare our method with Sound Morphing Toolbox (SMT) (Caetano, 2019), which is a set of Matlab functions targeting on musical instrument morphing that implement a sound morphing algorithm for isolated musical instrument sounds. Since SMT does not offer guidance for selecting perceptually uniform morph factors, we uniformly interpolate 11 morph factors in $[0,1]$.
Results and analysis. We set $N=11$ for SoundMorpher with an initial prompt ‘a music composition by {instrument}’. The comparison of our method and the baseline on timbral morphing is in Table 1. Overall, SoundMorpher demonstrates superior morphing quality compared to STM across various metrics, including audio quality, intermediateness, smoothness, and resynthesis quality   [^2] Since we perform timbral morphing within the same music composition, MFCCsE may not a suitable metric under the same musical content. In contrast, we focus on evaluating smoothness and intermediateness., when applied to different types of musical instrument timbre morphing. Notably, STM fails in Taiko-Hihat timbral morphing due to significant high reconstruction perceptual error. In contrast, SoundMorpher maintains robustness across different types of musical instruments, making it a more flexible and efficient solution for timbral morphing applications on different types of musical instruments. Figure. 1 provides a visualization of normalized timbre space, illustrating morphing trajectories generated by SMT and SoundMorpher. The timbre space is defined by three important timbral features: Log-Attack Time, Spectral Centroid, and Spectral Flux (McAdams et al., 1995; McAdams, 2013). The SMT trajectory shows distinct steps, indicating that the transitions between each intermediate sound are relatively abrupt. The spacing between the blue points suggests that each step represents a significant change in timbre, which may result in a less smooth perceptual transition between two musical instruments. In contrast, the trajectory produced by SoundMopher demonstrates a smoother curve. The points are more closely spaced, indicating more gradual changes between each intermediate timbre. This suggests that SoundMopher achieves a more continuous and natural-sounding morphing process, with each step being a smaller, more refined adjustment compared to SMT. Figure 7 and Figure 8 in the appendix provides additional visualization for this experiment.
5.3 Environmental Sound Morphing
Environmental sounds are used in video game production to provide a sense of presence within a scene. For example, in video, AR and VR games, sound morphing could enhance user immersion by adapting audio cues to specific visual and interactive contexts. This means that it could be useful to morph between sonic locations, e.g., a city and a park, or between sound effects, e.g., different animal sounds to represent fantasy creatures. In this experiment, we perform cyclostationary morphing with $N=5$ by SoundMorpher across various types of environmental sounds.
Dataset. We use ESC50 (Piczak, 2015) which consists of 5-second recordings organized into 50 semantic classes which loosely arranged into 5 major categories. We randomly select 4 major categories of scenarios to verify our method, including (1) Dog-Cat (animals voices), (2) Laughing-Crying baby (human sounds), (3) Church bells-Clock alarm (urban noise-interior sound), (4) Door knock-clapping (interior sounds-human sounds). Each category of scenarios contains 25 randomly selected audio pairs, thereby, 100 randomly paired samples in total.
Table 2: Environmental sound morphing with different types of environmental sounds.
[…]
Dog $\leftrightarrow$ Cat 26.08 0.081 17.77 73.92 1.293 0.323 $\pm$ 0.160 0.236
Laughing $\leftrightarrow$ Crying baby 10.39 0.044 9.35 65.98 0.855 0.214 $\pm$ 0.077 0.289
Church bells $\leftrightarrow$ Clock alarm 68.29 0.058 22.89 75.77 2.205 0.551 $\pm$ 0.299 0.312
Door knock $\leftrightarrow$ Clapping 21.36 0.083 10.85 76.35 1.594 0.428 $\pm$ 0.220 0.321
[…]
5.4 Music Morphing
Film or game post-production often requires blending or fading between music tracks to seamlessly transition background music in between scenes. Music morphing transitions between two music compositions without cross fading, that is, each moment of the morphed music would be a single composition with elements that are perceptually in between both source and target music, rather than simply blending the two together. Different from timbral morphing, music morphing could ideally be accomplished with compositions from different genres and mixed musical instruments. In this experiment, we use SoundMorpher to perform dynamic morphing on music with $N=15$.
Dataset. In this experiment, we randomly selected 50 sample pairs from 20 musical samples available on AudioLDM2 (Liu et al., 2024) demonstration page. These 10-second music compositions that span different genres and feature both single or mixed musical instrument arrangements.
Figure 2: Visualization of dynamic morphed music with $N=15$, source music and target music.
Results and analysis. In this experiment, we select an initial text prompt as ‘a sound clip of music composition’ to perform dynamic morphing. Even though this experiment contains morph complex music compositions with different music genres and music instruments, Table 3 shows our method still superiors on perceptual smoothly transiting source music to the target music and ensures correspondence, intermediateness and smoothness. Figure 2 provides strong visual evidence that the dynamic morphing method effectively transitions from the source to the target music while maintaining perceptual smoothness, correspondence, and intermediate transformations. The spectrogram illustrates that the morphed music transitions gradually, maintaining smooth spectral changes over time, which suggests the method successfully morphs the source into the target music.
Table 3: Music morphing experimental results & ablation study for sound perceptual features, where N represents the number of components of PCA for reducing dimension of Mel-spectrogram.
[…]
Reduced Mel-Spec. (N=2) 0.187 10.31 58.57 0.793 0.056 $\pm$ 0.075 0.182
Reduced Mel-Spec. (N=3) 0.151 10.76 59.39 0.779 0.055 $\pm$ 0.079 0.151
Mel-Spec. 0.056 9.85 56.09 0.847 0.068 $\pm$ 0.045 0.178
MFCCs 0.053 10.11 57.38 0.987 0.071 $\pm$ 0.050 0.156
Spectral contras 0.066 10.54 58.44 0.863 0.061 $\pm$ 0.071 0.155
Table 4: Mean opinion score study on environmental sound morphing and music morphing task.
[…]
Environmental sound morphing 3.78 $\pm$ 0.31 3.67 $\pm$ 0.40 3.57 $\pm$ 0.47 3.67 $\pm$ 0.39
Music morphing 3.81 $\pm$ 0.39 3.55 $\pm$ 0.51 3.49 $\pm$ 0.48 3.62 $\pm$ 0.46
5.5 Discussion
Mean opinion score (MOS) study. We conducted a MOS study as a subjective evaluation for the correspondence, intermediateness, and smoothness of morphed results from SoundMorpher. The study involved 21 volunteers, and detailed methodology is in Appendix 9. As shown in Table 4, the results suggest that SoundMorpher is versatile, performing similarly well across both music and environmental sound morphing tasks, with no significant differences observed in the overall MOS. This consistency in scores indicates that SoundMorpher effectively handles the unique challenges posed by the distinct characteristics of music and environmental sounds, such as the continuous nature of music compared to the more segmented structure of environmental sounds. Despite the objective metric results showing clear differences between the two tasks, the human evaluation suggests that SoundMorpher remains robust across different sound types. One possible interpretation is that our objective metrics are more sensitive to variations in the measured aspects than participants.
Table 5: Comparison with MorphFader
Method CDPAMT CDPAMmean±std MFCCsE
MorphFader 0.972 0.243 $\pm$ 0.139 0.065
SoundMorpher 0.935 0.226 $\pm$ 0.162 0.065
Model comparison. We compare SoundMorpher with a concurrent work, MorphFader (Kamath et al., 2024), based on the criteria outlined in Section 3.1. MorphFader relies on a pre-trained AudioLDM (Liu et al., 2023) and perform sound morphing by text instructions. We compare with 7 examples provided on the demonstration of MorphFader, that MorphFader uniformly samples $5$ morph factors in $[0,1]$, resulting in a morph path with $[0,0.25,0.5,0.75,1]$. In contrast, SoundMorpher finds $\alpha$ values according to constant $\Delta p$ with $5$ uniformly interpolate $p$ points by binary search. As in Table 5, SoundMorpher produces a smoother and perceptual intermediates morphing than MorphFader. See Appendix 10 for details.
Ablation study on sound perceptual features. We verified the perceptual feature in SPDP in music morphing task. We select alternative music information retrieval features (MIR) including MFCCs with 13 coefficients (Logan et al., 2000), and spectral contrast (Jiang et al., 2002). We use principal component analysis (PCA) to reduce the dimensionality of Mel-spectrogram to further capture variation of spectral content over time, which is referred to as reduced Mel-Spec. (Stevens et al., 1937; Casey et al., 2008; Jiang et al., 2002). Table 3 shows performance comparisons of SoundMorpher with different features, SoundMorpher with Mel-spectrogram achieves better morphing quality in terms of correspondence and smoothness variation with smaller FAD, FD and CDPAMstd. While Mel-spectrogram yields higher CDPAMmean, CDPAMT and MFCCsE compared to reduced Mel-Spec. and MFCCs, the differences in metric values are not significant. However, the overall morph quality with Mel-spectrogram is consistently better than other features. This suggests Mel-spectrogram, as a pseudo-3D representation, provide more perceptual and semantic information, which contributes to improve morph quality compared to higher-level features.
Table 6: Experiment and ablation study results on music morphing.
[…]
Informative $\checkmark$ 0.047 10.21 56.62 1.213 0.086 $\pm$ 0.069 0.166
Informative $\checkmark$ 0.044 10.21 56.13 1.077 0.084 $\pm$ 0.066 0.155
Uninformative $\checkmark$ 0.057 10.37 55.89 1.036 0.074 $\pm$ 0.049 0.211
Uninformative $\checkmark$ 0.056 9.85 56.09 0.847 0.068 $\pm$ 0.045 0.178
Uninformative v.s. informative initial text prompt. Complex audio usually cannot easily yield precise information to users. For example, it is a challenge for non-professional users to describe the genre of a music. We conduct an ablation study for initial text prompt on music morphing to verify effectiveness of text-guided conditional embedding optimization. We use a general initial text prompt, ‘a sound clip of music composition.’, as an uninformative initial prompt. And we use the given text prompts in AudioLDM2  [^3] https://audioldm.github.io/audioldm2/ as informative inital prompts. As in Table 6, informative initial text prompts may help with resynthesis quality and further improves morph correspondence. Despite the improved resynthesis quality with informative initial text prompts, the results show a decline in morphing intermediateness and smoothness. One possible reason is the better resynthesis quality makes the resynthesis endpoints more distinct (i.e., larger semantic gap), which could lead to slight decline in intermediateness and smoothness. However, the performance difference on initial text prompts is not significant which illustrates effectiveness of conditional embedding optimization.
Inference steps. In our experiment, we follow the configuration of Zhang et al. (2024) and set DDIM steps to 100. To verify whether DDIM steps affect SoundMorpher performances, we compare with 20 DDIM steps in Table 6. Larger inference step seems to help for reconstruction quality and slightly imporves morph quality, however, performance differences between inference steps are not significant. This indicates SoundMorpher is robust for inference steps, and we extend this ablation study on environmental sound morphing task in Appendix 11.1. Thus, we suggest selecting a suitable DDIM step to trade-off overall binary search algorithm time-consuming and morph quality.
Limitations. The current implementation of SoundMorpher based on AudioLDM2 with 16.0kHz sampling rate, which may limit output audio quality. The conditional embeddings optimization only applies to sounds that can be produced by AudioLDM2. Sound examples that close to white noise, such as pure wind blowing used in MorphGAN (Gupta et al., 2023) are not easily generated by AudioLDM2, which makes the conditional embedding optimization produce low quality resynthesis sounds. We also observed that input sounds with a large semantic gap (e.g., Church bells-Clock alarm in Table 2) result in lower morphing quality. Furthermore, we observed when two audios exhibit significant temporal structure differences, such as environmental sounds, SoundMorpher may produce abrupt transitions, see Appendix 11.4 and Figure 6 for further details.
6 Conclusion
We propose SoundMorpher, a sound morphing method base on a pretrained diffusion model that produces perceptually uniform morphing trajectories. Unlike existing methods, we refined the sound morphing problem and explored an explicit connection between morph factor and perceptual stimuli of morphed results which offers better flexibility and higher morphing quality, making it adaptable to various morphing methods and real-world scenarios. We validate SoundMorpher by a series of objective quantitative metrics as well as mean opinion score study following criteria proposed by Caetano & Osaka (2012). These quantitative objective metrics may help to formalize future studies on sound morphing evaluation. Furthermore, we demonstrated that SoundMorpher can be applied to wide range of real-world applications in our experiments and conducted in-depth discussions. SoundMorpher also has the potential to achieve voice morphing, as its foundational model AudioLDM2 supports speech generation; however, we leave this exploration for future work.
Appendix
7 SoundMorper implementation details
This section provides further details on SoundMorpher pipeline and implementations.
7.1 Overall pipeline of SoundMorpher
The overall pseudo pipeline for SoundMorpher is provided in Algorithm 1, where the overall pipeline of SoundMorpher contains three parts: (1) Conditional embedding optimization. (2) Model adaptation. (3) perceptually uniform binary search with constant SPDP increment.
Algorithm 1 Pipeline of SoundMorpher
A pre-trained AudioLDM2 pipeline including a pre-trained VAE with a encoder $g_{\theta}$ and a decoder $g_{\phi}$, a pre-trained latent diffusion model $\epsilon_{\theta}$, a pre-trained T5 model $f_{\phi}$, and a pre-trained GPT-2 model $f_{\varphi}$. Learning rates $\eta_{1},\eta_{2}$. Source and target audios $x^{(0)}$ and $x^{(1)}$. An initial text prompts $y$. perceptually uniform interpolation number $N$. Tolerance error for binary search $\epsilon_{tolerance}$. Number of training steps for text inversion for conditional embedding optimization $T_{inv}$. Number of training steps for model adaptation $T_{adapt}$. LoRA rank $r$, Number of steps for DDIM $T$.
Start morph factor $\alpha_{start}=0$, end morph factor $\alpha_{end}=1$. Start perceptual point $p_{start}=[0,1]$, and end perceptual point $p_{end}=[1,0]$.
Initialize: $z_{0}^{(0)}=g_{\theta}(x_{0}^{(0)})$, $z_{0}^{(1)}=g_{\theta}(x_{0}^{(1)})$; $E^{0}=[f_{\phi}(y),f_{\varphi}(y)]$, $E^{1}=[f_{\phi}(y),f_{\varphi}(y)]$;
# Step 1: Text-guided conditional embedding optimization.
for $i$ from $1$ to $T_{inv}$ do
     Randomly sample time step $t$ and random noise $\epsilon_{t}\sim N(0,I)$.
     Adding noise to data $z_{t}^{(0)}\leftarrow\sqrt{\gamma_{t}}z_{0}^{(0)}+\sqrt{(1-\gamma_{t})}\epsilon_{t},z_{t}^{(1)}\leftarrow\sqrt{\gamma_{t}}z_{0}^{(1)}+\sqrt{(1-\gamma_{t})}\epsilon_{t}$.
     $E^{(0)}\leftarrow E^{(0)}-\eta_{1}\nabla_{E^{(0)}}\mathcal{L}_{DPM}(z_{0}^{(0)},E^{(0)};\theta)$.
     $E^{(1)}\leftarrow E^{(1)}-\eta_{1}\nabla_{E^{(1)}}\mathcal{L}_{DPM}(z_{0}^{(1)},E^{(1)};\theta)$.
end for
# Step 2: Model adaptation with LoRA.
for $i$ from $1$ to $T_{adapt}$ do
     Model adaptation with LoRA according to Equation 9 and Equation 10 with $\eta_{2}$ learning rate.
end for
# Step 3: Perceptual-uniform binary search with constant SPDP increment.
Obtaining initial latent states $z_{T}^{(0)}$ and $z_{T}^{(1)}$ by Equation 4
$p_{list}\leftarrow$ ConstantSPDP$(N,p_{start},p_{end})$triangleright Obtain target SPDP points by Algorithm 2
$\alpha_{list}\leftarrow\text{BinarySeach}(\alpha_{start},\alpha_{end},p_{list},\epsilon_{tolerance})$
for $\alpha$ in $\alpha_{list}$ do
     $E^{(\alpha)}\leftarrow(1-\alpha)E^{(0)}+\alpha E^{(1)}$
     $z_{T}^{(\alpha)}\leftarrow\frac{sin(1-\alpha)w}{sinw}z_{T}^{(0)}+\frac{sin\alpha w}{sinw}z_{T}^{(1)}$
     for $t$ from T todo
         $z_{t-1}^{(\alpha)}\leftarrow\sqrt{\gamma_{t-1}}(\frac{z_{t}-\sqrt{1-\gamma_{t}}\hat{\epsilon}_{\theta}^{(t)}(z_{t},E^{(\alpha)})}{\sqrt{\gamma_{t}}})+\sqrt{1-\gamma_{t-1}}\hat{\epsilon}_{\theta}^{(t)}(z_{t},E^{(\alpha)})$
     end for
end for
$x^{(\alpha)}\leftarrow\text{vocoder}(g_{\phi}(z_{0}^{(\alpha)}))$triangleright Decode latent variable and obtain audio waveform by a vocoder.
return $\{x^{(\alpha)}\}_{\alpha\in\alpha_{list}}$
7.2 Model adaptation with LoRA
In the task of image morphing, Yang et al. (2023) indicate adapting the model to the input pair helps to limit the degree of morphed variation by suppressing high-density regions that are not related to the given images. Compared to vanilla fine-tuning approaches, LoRA has advantages in training efficiency with injecting limited trainable parameters. The model adaptation can be defined as
$\text{min}_{\Delta\theta^{\prime}}\mathcal{L}_{DPM}(z_{0}^{(0)},E^{(0)};\theta+\Delta\theta^{\prime})+\text{min}_{\Delta\theta^{\prime}}\mathcal{L}_{DPM}(z_{0}^{(1)},E^{(1)};\theta+\Delta\theta^{\prime})\text{ s.t. rank}(\Delta\theta^{\prime})=r$ (9)
where $r$ represents LoRA rank.
Unconditional Bias Correction. To achieve high text alignment during inference time, we use additional LoRA parameters $\Delta\theta_{0}$ with a small rank $r_{0}$ to perform bias correction, as
$\text{min}_{\Delta\theta_{0}}\mathcal{L}_{DPM}(z_{0}^{(0)},\varnothing;\theta+\Delta\theta_{0})+\text{min}_{\Delta\theta_{0}}\mathcal{L}_{DPM}(z_{0}^{(1)},\varnothing;\theta+\Delta\theta_{0})\text{ s.t. rank}(\Delta\theta_{0})=r_{0}$ (10)
During inference, with $\theta^{\prime}=\theta+\Delta\theta^{\prime}$ and $\theta_{0}=\theta+\Delta\theta_{0}$, the noise prediction becomes
$\hat{\epsilon}_{\theta}(z_{t},t,E):=w\epsilon_{\theta^{\prime}}(z_{t},t,E)+(1-w)\epsilon_{\theta_{0}}(z_{t},t,\varnothing).$ (11)
In our experiment, we set $r=4$ and $r_{0}=2$. Although Yang et al. (2023) provide a heuristic suggestion for setting the LoRA rank value for image morphing task, however, we further investigate the relationship between LoRA rank $r$ and method performance in Table 8 in sound morphing task and discussion in Appendix 11.2.
7.3 perceptually uniform binary search with constant SPDP increment
Algorithm 2 provides detail pesudocodes for how to implement perceptually uniform binary search with constant SPDP increment. This pseudo algorithm includes two steps, firstly, compute constant SPDP increment according to interpolte point number $N$ and obtain $N$ target SPDP points as $\{p_{i}\}_{i=1}^{N}$. Secondly, perform binary search to find correponding morph factor $\alpha$ series $\{\alpha_{i}\}_{i=1}^{N}$ according to $\{p_{i}\}_{i=1}^{N}$.
Algorithm 2 Pseudo algorithm for perceptually uniform binary search with constant SPDP increment
$\alpha_{start}$: starting alpha value; $\alpha_{end}$: ending alpha value; $N$: number of interpolations; source audio $x^{(0)}$; target audio $x^{(1)}$;
$p_{list}=[]$; $p_{start}=[0,1]^{T}$; $p_{end}=[1,0]^{T}$; $\alpha_{list}=[]$;
# Step 1: Obtain target SPDP points with constant increment.
Procedure ConstantSPDP$(N,p_{start},p_{end})$
for $i$ from $1$ to $N-1$ do $\triangleright$ Find target SPDP points with constant $\Delta p$.
     $t\leftarrow\frac{i}{N-1}$;
     $p_{i}\leftarrow(1-t)\times p_{start}+t\times p_{end}$
     $p_{list}\leftarrow p_{list}\cup[p_{i}]$;
end for
# Step 2: Perform binary search given target SPDP points with constant $\Delta p$.
Procedure BinarySearch($\alpha_{start},\alpha_{end},x^{(0)},x^{(1)},p_{list},\epsilon_{tolerance}$);
$\alpha_{list}\leftarrow[\alpha_{start}]$;
$\alpha_{cur}\leftarrow\alpha_{start}$;
for $p_{i}$ from $p_{1}$ to $p_{N-2}$ do
     $p_{target}\leftarrow p_{i}$
     $\alpha_{t1}\leftarrow\alpha_{cur}$;
     $\alpha_{t2}\leftarrow\alpha_{end}$;
     $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$;
     $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$triangleright Compute SPDP by Equation 7.
     while $|p_{mid}-p_{i}|>\epsilon_{tolerance}$ do $\triangleright$ Perform binary search
         if $p_{mid}>p_{target}$ then
              $\alpha_{t2}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
         else
              $\alpha_{t1}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
         end if
         $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
         $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$
     end while
     $\alpha_{list}\leftarrow\alpha_{list}\cup[\alpha_{mid}]$triangleright Append the reult to $\alpha_{list}$.
end forreturn $\alpha_{list}$;
Algorithm 3 Pseudo algorithm for static morphing
Source audio $x^{(0)}$, target audio $x^{(1)}$, specific SPDP point $p$, tolerance error for binary search $\epsilon_{tol}$, number of steps for DDIM $T$, VAE decoder $g_{\phi}$.
$\alpha_{start}=0,\alpha_{end}=1$<

### v2

@@
 Abstract
-We present SoundMorpher, a sound morphing method that generates perceptually uniform morphing trajectories using a diffusion model.
-Traditional sound morphing methods models the intractable relationship between morph factor and perception of the stimuli for resulting sounds under a linear assumption, which oversimplifies the complex nature of sound perception and limits their morph quality.
-In contrast, SoundMorpher explores an explicit proportional mapping between the morph factor and the perceptual stimuli of morphed sounds based on Mel-spectrogram.
-This approach enables smoother transitions between intermediate sounds and ensures perceptually consistent transformations, which can be easily extended to diverse sound morphing tasks.
-Furthermore, we present a set of quantitative metrics to comprehensively assess sound morphing systems based on three objective criteria, namely, correspondence, perceptual intermediateness, and smoothness.
-We provide extensive experiments to demonstrate the effectiveness and versatility of SoundMorpher in real-world scenarios, highlighting its potential impact on various applications such as creative music composition, film post-production and interactive audio technologies.
-See SoundMorpher-demo for the listening demonstration.
-1 Introduction
+We present SoundMorpher, an open-world sound morphing method designed to generate perceptually uniform morphing trajectories.
+Traditional sound morphing techniques typically assume a linear relationship between the morphing factor and sound perception, achieving smooth transitions by linearly interpolating the semantic features of source and target sounds while gradually adjusting the morphing factor.
+However, these methods oversimplify the complexities of sound perception, resulting in limitations in morphing quality.
+In contrast, SoundMorpher explores an explicit relationship between the morphing factor and the perception of morphed sounds, leveraging log Mel-spectrogram features.
+This approach further refines the morphing sequence by ensuring a constant target perceptual difference for each transition and determining the corresponding morphing factors using binary search.
+To address the lack of a formal quantitative evaluation framework for sound morphing, we propose a set of metrics based on three established objective criteria.
+These metrics enable comprehensive assessment of morphed results and facilitate direct comparisons between methods, fostering advancements in sound morphing research.
+Extensive experiments demonstrate the effectiveness and versatility of SoundMorpher in real-world scenarios, showcasing its potential in applications such as creative music composition, film post-production, and interactive audio technologies.
+Index Terms:
+Sound morphing, open-world, diffusion models.
+publicationid: pubid: $0$–$00$ 2021 IEEE
+I Introduction
 Sound morphing is a technique to create a seamless transformation between multiple sound recordings.
-The goal is to produce perceptual intermediate sounds that gradually change from one sound to another.
-Sound morphing has a wide range of applications, including music compositions, synthesizers, psychoacoustic experiments to study timbre spaces (Caetano & Rodet, 2011; Hyrkas, 2021), and practical applications such as film post-production, AR or VR interactive games, and adaptive audio content in video games (Qamar et al., 2020; Siddiq, 2015).
-Traditional sound morphing methods used the interpolation principle of sound synthesis technique, which relies on interpolating the parameters of a sinusoidal model (Tellman et al., 1995; Osaka, 1995; Williams et al., 2014).
-Others make use of digital signal processing techniques to explore high-level audio features in the time-frequency domain to achieve more effective and continuous morphing (Williams et al., 2014; Brookes & Williams, 2010; Caetano & Rodet, 2010; 2011; Roma et al., 2020; Caetano, 2019).
-However, these methods are limited to applications such as producing inharmonic and noisy environmental sounds (Gupta et al., 2023; Kamath et al., 2024).
-Despite the increasing interest in applying machine learning to sound generation, there has only been limited exploration in sound morphing.
-Recent approaches (Zou et al., 2021; Gupta et al., 2023; Kim et al., 2019b; Kamath et al., 2024) have shown their superior effectiveness compared to traditional methods in various scenarios.
-However, we observed several critical limitations of those existing methods.
-Firstly, they are primarily designed for static or cyclostationary morphing (see Sec.
-3.1), limiting their applicability to dynamic sound transformations.
-Secondly, these approaches often lack sufficient quantitative evaluation, limiting further analysis of their effectiveness.
-Thirdly, they require training on task-specific datasets, which limits their application in different scenarios.
-Most importantly, they typically assume a linear relationship between morphing factors and sound perception, and achieve smooth morphing by gradually changing the morph factor.
+The goal is to produce intermediate sounds that might exist between two sounds and contrasts with simpler mixing or cross-fading between sounds.
+Sound morphing has a wide range of applications, including music compositions, synthesizers, psychoacoustic experiments to study timbre spaces [1, 2], film post-production, AR/VR, and adaptive audio content in video games [3, 4].
+Traditionally, sound morphing relied on interpolating the parameters of a sinusoidal synthesis model [5, 6, 7].
+High-level audio features in the time-frequency domain can also be controlled to achieve more effective and continuous morphing [7, 8, 9, 1, 10, 11].
+These techniques have quality limitations that rule out applications in inharmonic and noisy sounds such as environmental sounds [12, 13].
+Increasing interest in sound generation using machine learning (ML) has delivered techniques for ML-based sound morphing [14, 12, 15, 13] that out-perform traditional methods in scenarios such as timbre morphing, audio texture morphing and environmental sound morphing; however, we observed critical limitations of these existing sound morphing methods.
+Firstly, they are primarily designed for specific morphing methods or tailored to particular application scenarios on task-specific datasets, which limits their ability to broader applications and different scenarios.
+Secondly, a lack of sufficient and formal quantitative evaluation limits further analysis of their effectiveness [12] and impedes fair comparisons with other sound morphing techniques.
+Most importantly, these methods typically assume a linear relationship between morphing factors and sound perception, as proposed by [16].
+They achieve smooth morphing by interpolating semantic features linearly between the source and target sounds and gradually varying the morphing factor.
 This assumption oversimplifies the complex nature of sound perception, as gradually changing morph factors does not inherently result in smooth perceptual transitions.
-To this end, our goal is to develop a method that achieves perceptually coherent morphing, ensuring seamless and natural sound transition.
-In this paper, we introduce SoundMorpher, a sound morphing method that produces perceptually smooth and intermediate morphing, comprising the following key contributions.
-  • SoundMopher is the first open-world sound morphing method based on a pre-trained diffusion model, which integrates typical morph tasks such as static, dynamic and cyclostationary morphing.
-Unlike prior works (Kim et al., 2019b; Gupta et al., 2023), SoundMorpher can be broadly applied to various real-world tasks without requiring extensive retraining.
-  • We propose the sound perceptual distance proportion (SPDP), which explicitly connects morph factors and perceptual stimuli of morphed results.
-This allows SoundMorpher to produce morphing paths with a uniform change in perceptual stimuli, achieving more seamless perceptual transitions compared to existing methods (Kamath et al., 2024).
-  • We adapt a set of comprehensive quantitative metrics according to criteria proposed by Caetano & Osaka (2012) for evaluation, addressing the lack of quantitative assessment for sound morphing systems (Caetano, 2019; Zou et al., 2021; Caetano & Rodet, 2013) and may offer insights for analyzing and comparing future sound morphing methods.
-  • We provide extensive experiments to demonstrate that SoundMorpher can be effectively applied to several potential applications in broader real-world scenarios, including musical instrument timbre morphing, music morphing and environmental sound morphing.
-2 Related Work
-In this section, we first present a detailed review of related works on sound morphing task. Then, we also briefly introduce tasks that are similar to sound morphing and clarify the differences.
+Our goal is to address these limitations by introducing SoundMorpher, a flexible sound morphing method, which can be broadly applied to different sound morphing tasks to achieve perceptually coherent morph, ensuring seamless and natural sound transition.
+In this work we:
+  • Introduce the first open-world sound morphing method based on a pre-trained diffusion model, which integrates typical morph tasks such as static, dynamic and cyclostationary morphing.
+Unlike [15, 12], SoundMorpher can be broadly applied to various real-world sound morphing tasks without requiring extensive retraining on an additional dataset, paving the way for future sound morphing studies to focus on leveraging pre-trained audio generation models for greater efficiency and versatility.
+  • Propose sound perceptual distance proportion (SPDP), which explicitly connects morph factors and perception of morphed results.
+This allows SoundMorpher to produce morphing paths with a uniform change in perceptual stimuli, achieving more seamless perceptual transitions compared to existing methods [13].
+  • Adapt established criteria [16] for quantitative evaluation, addressing the lack of comprehensive objective quantitative assessment for sound morphing systems [11, 14, 17].
+  • Provide extensive experiments to demonstrate that SoundMorpher can be effectively applied to various applications in real-world scenarios including timbre morphing, music morphing and environmental sound morphing.
+II Related Work
+In this section, we present a detailed review of related works on sound morphing task and briefly introduce contrasts with other similar tasks.
 Sound morphing.
-Traditional sound morphing methods rely on interpolating parameters of a sinusoidal sound synthesis model (Tellman et al., 1995; Osaka, 1995; Williams et al., 2014; Primavera et al., 2012).
-To achieve more effective and continuous morphing, Williams et al.
-(2014); Brookes & Williams (2010); Caetano & Rodet (2010; 2011); Roma et al.
-(2020); Caetano (2011) target on exploring perceptual spectral domain audio features by digital signal processing techniques, such as MFCCs, spectral envelope, etc..
-Others such as Kazazis et al.
-(2016) involve a hybrid approach that extracts audio descriptors to morph accordingly and interpolate between the spectrotemporal fine structures of two endpoints according to morph factors.
-Machine learning sound morphing methods offer advantages such as high morphing quality by leveraging semantic representation interpolation within a model instead of traditional audio feature interpolation.
-Zou et al.
-(2021) proposes a non-parallel many-to-one static timbre morphing framework that integrates and fine-tunes the machine learning technique (i.e., DDSP-autoencoder (Engel et al., 2020)) with spectral feature interpolation (Caetano & Rodet, 2013).
-Kim et al.
-(2019b) targets synthesizing music corresponding to a note sequence and timbre, which uses non-linear instrument embedding as timbre control parameters under a pretrained WaveNet (Engel et al., 2017) to achieve timbre morphing between instruments.
-Luo et al.
-(2019) learns latent distributions of VAEs to disentangle representations for pitch and timbre of musical instrument sounds.
-Tan et al.
-(2020) uses a GM-VAE to achieve style morphing to generate realistic piano performances in the audio domain following temporal style conditions for piano performances, which morphs the conditions such as onset roll and MIDI note into input audio.
-MorphGAN (Gupta et al., 2023) targets on audio texture morphing by interpolating within conditional parameters, and trained the model on a water-wind texture dataset.
-A recent concurrent work by Kamath et al.
-(2024) uses a pre-trained AudioLDM (Liu et al., 2023) to morph sound between two text prompts.
-In contrast, we focus on classical sound morphing, where the morphing process is performed directly between two given audios rather than between text prompts.
-A key advantage of our method is its ability to provide precise guidance during the morphing process, as the target audio delivers exact information on how the source sound should evolve—something that text prompts cannot always achieve, for example, morphing between two music compositions.
+Traditional sound morphing methods are based on interpolating parameters of a sinusoidal sound synthesis model [5, 6, 7, 18].
+To achieve more effective and continuous morphing, [7, 8, 9, 1, 10, 19] exploring perceptual spectral domain audio features by digital signal processing techniques, such as MFCCs, spectral envelope, etc.
+A hybrid approach has been explored that extracts audio descriptors to morph accordingly and interpolates between the spectrotemporal fine structures of two endpoints according to morph factors [20].
+Machine learning sound morphing methods offer advantages such as high morphing quality by interpolation of semantic representations within a model instead of traditional audio features.
+[14] proposes a non-parallel many-to-one static timbre morphing framework that integrates and fine-tunes the machine learning technique (i.e., DDSP-autoencoder [21]) with spectral feature interpolation [17].
+[15] synthesizes music corresponding to a note sequence and timbre, using nonlinear instrument embeddings as timbre control parameters under a pretrained WaveNet [22] to achieve timbre morphing between instruments.
+[23] learns latent distributions of VAEs to disentangle representations of the pitch and timbre of musical instrument sounds.
+[24] uses a GM-VAE to achieve style morphing to generate realistic piano performances in the audio domain following temporal style conditions for piano performances, which morphs the conditions such as onset roll and MIDI note into input audio.
+MorphGAN [12] targets audio texture morphing by interpolating within conditional parameters training the model on a water-wind texture dataset.
+A recent concurrent work [13] uses a pre-trained AudioLDM [25] to morph sound between two text prompts.
+In contrast, we focus on classical sound morphing, where the morphing process is performed directly between two given sounds rather than between text prompts.
+A key advantage of our method is its ability to provide precise guidance during the morphing process, since the target audio delivers exact information on how the source sound should evolve—something that text prompts cannot always achieve, for example, morphing between two music compositions.
 Synthesizer preset interpolation.
-Synthesizer preset interpolation achieves sound morphing by developing models that compute interpolations within the domain of synthesis parameters for a black-box synthesizer (Le Vaillant & Dutoit, 2023; Dutoit et al., 2023; Le Vaillant & Dutoit, 2024).
+A form of sound morphing can be achieved by developing models that compute interpolations in the parameters for a black-box synthesizer [26, 27, 28].
 Unlike classical sound morphing, which perceptually blends two audio files into an intermediate sound, synthesizer preset interpolation treats the synthesizer as a non-differentiable black box, with presets composed of both numerical and categorical parameters.
@@
 Text-to-audio editing is the process of using text queries to edit audio.
-With the success of diffusion models in image editing tasks, recent works target zero-shot audio editing with text instructions (Manor & Michaeli, 2024; Zhang et al., 2024; Lan et al., 2024) involving tasks such as inpainting, outpainting, timbre transfer, music genre transfer, or vocals removal.
+With the success of diffusion models in image editing tasks, recent works target zero-shot audio editing with text instructions [29, 30, 31] involving tasks such as inpainting, outpainting, timbre transfer, music genre transfer, or vocals removal.
 Timbre transfer.
 Timbre transfer is a specific task that aims at converting the sound of a musical piece by one instrument (source) into the same piece played by another instrument (target).
-This concerns the task of converting a musical piece from one timbre to another while preserving the other music-related characteristics (Comanducci et al., 2024; Jain et al., 2020; Li et al., 2024).
+This concerns the task of converting a musical piece from one timbre to another while preserving the other music-related characteristics [32, 33, 34].
 Voice conversion and morphing.
-Voice conversion (VC) involves modifying vocal characteristics of a source speech to match a target speaker, either by using target speeches or text (Li et al., 2023; Yao et al., 2024; Niu et al., 2024; Sheng et al., 2024).
-The primary objective of VC is to alter the vocal identity to closely resemble the target voice style, while preserving the linguistic content of the source speech.
+Voice conversion (VC) involves modifying vocal characteristics of a source speech to match a target speaker, either by using target speeches or text [35, 36, 37, 38].
+The primary objective of VC is to alter the vocal identity to closely resemble the target voice style while preserving the linguistic content of the source speech.
 Voice morphing is a broader scope, focusing on blending or transforming one voice into another.
-This often involves creating an intermediate voice that incorporates characteristics of both source and target voices, allowing for gradual transitions between them (Sheng et al., 2024).
-3 Preliminaries
-3.1 Sound Morphing
-Sound morphing aims to produce intermediate sounds as different combinations of model source sound $\hat{S}_{1}$ and target sound $\hat{S}_{2}$ (Caetano & Rodet, 2010; 2011), which can be formulated as
+This often involves creating an intermediate voice that incorporates characteristics of both source and target voices, allowing for gradual transitions between them [38].
+III Preliminaries
+III-A Sound Morphing
+Sound morphing aims to produce intermediate sounds as different combinations of model source sound $\hat{S}_{1}$ and target sound $\hat{S}_{2}$ [9, 1], which can be formulated as
 $M(\alpha,t)=(1-\alpha(t))\hat{S}_{1}+\alpha(t)\hat{S}_{2}$ (1)
-Each step is characterized by one value of a single parameter $\alpha$, the so-called morph factor, which ranges between 00 and $1$, where $\alpha=0$ and $\alpha=1$ produce resynthesized source and target sounds, respectively.
-Due to the intrinsic temporal nature of sounds, sound morphing usually involve three main types: dynamic morphing, where $\alpha$ gradually transfers from 00 to $1$ over time dimension (Kazazis et al., 2016), static morphing, where a single morph factor $\alpha$ leads to an intermediate sound between source and target (Sethares & Bucklew, 2015), and cyclostationary morphing where several hybrid sounds are produced in different intermediate points (Slaney et al., 1996).
-To solve the limitation on previous works that target on expensive perceptual evaluation only, Caetano & Osaka (2012) proposes three objective criteria for sound morphing techniques: (1) Correspondence.
+Each step is characterized by one value of a single parameter $\alpha$, the so-called morph factor, which ranges between 00 and $1$.
+$\alpha=0$ and $\alpha=1$ produce resynthesized source and target sounds, respectively.
+Due to the intrinsic temporal nature of sounds, sound morphing usually involves three main types: dynamic morphing, where $\alpha$ gradually transfers from 00 to $1$ over time [20], static morphing, where a single morph factor $\alpha$ leads to an intermediate sound between source and target [39], and cyclostationary morphing where several hybrid sounds are produced in different intermediate points [40].
+To solve the limitation on previous works that target on expensive perceptual evaluation only, [16] provided three objective criterions to formalize the evaluation for sound morphing methods: (1) Correspondence.
 The morph is achieved by a description whose elements are intermediate between source and target sounds, highlighting semantic level transition; (2) Intermediateness.
 The morphed objects should be perceived as intermediate between source and target sounds, evaluating perceptual level correlation; (3) Smoothness.
-The morphed sounds should change gradually (i.e., ‘smoothly’) from source to target sounds, by the same amount of perception increment.
-Under the assumption of linear perceptual stimuli, adding the same factor should increase the same amount of perception.
-In this study, we evaluate SoundMorpher according to the three criteria by a series of comprehensive objective quantitative metrics in addition to perceptual evaluation.
-3.2 Latent Diffusion Model on Audio Generation
-SoundMorpher utilizes a pretrained text-to-audio (TTA) latent diffusion model (LDM) (Rombach et al., 2022) to achieve sound morphing.
+The morphed sounds should change gradually from source to target sounds, by the same amount of perception increment.
+Specifically, the linear assumption proposed by [16] is that adding the same amount of morph factor should increase the same amount of perception difference.
+In this study, we evaluate SoundMorpher according to the three criterions proposed by [16] with a series of comprehensive objective quantitative metrics.
+III-B Latent Diffusion Model on Audio Generation
+SoundMorpher uses a pretrained text-to-audio (TTA) latent diffusion model (LDM) [41] to achieve sound morphing.
 This approach offers the advantage of performing various types of sound morphing without the need to train the entire model or use additional datasets.
-Specifically, we use AudioLDM2 (Liu et al., 2024), a multi-modality conditions to audio model.
-It employs a pre-trained variational autoencoder (VAE) (Kingma & Welling, 2013) to compress audio $x$ into a low-dimension latent space as VAE representations $z$.
-AudioLDM2 generates latent variables $z_{0}$ from a Gaussian noise $z_{T}$ given the condition $C$ and further reconstruct audio $\hat{x}$ from $z_{0}$ by VAE decoder and a vocoder (Kong et al., 2020).
+Specifically, we use AudioLDM2 [42], a multi-modality conditions to audio model.
+It employs a pre-trained variational autoencoder (VAE) [43] to compress audio $x$ into a low-dimension latent space as VAE representations $z$.
+AudioLDM2 generates latent variables $z_{0}$ from a Gaussian noise $z_{T}$ given the condition $C$ and further reconstructs audio $\hat{x}$ from $z_{0}$ by VAE decoder and a vocoder [44].
 AudioLDM2 uses an intermediate feature $Y$ as an abstraction of audio data $x$ to bridge the gap between conditions $C$ and audio $x$, named language of audio (LOA).
-The LOA feature is obtained by a AudioMAE (Huang et al., 2022; Tan et al., 2024) and a series of post-processing formulated as $Y=\mathcal{A}(x)$.
+The LOA feature is obtained by a AudioMAE [45, 46] and a series of post-processing formulated as $Y=\mathcal{A}(x)$.
 The generation function $\mathcal{G}(\cdot)$ is achieved by a LDM.
-In the inference phase, AudioLDM2 approximates LOA feature by the given condition as $\hat{Y}=\mathcal{M}(C)$ using a fine-tuned GPT-2 model (Radford et al., 2019).
-Then generates audios conditioned on the estimated LOA feature $\hat{Y}$ and an extra text embedding $E_{T5}$ from a FLAN-T5 (Chung et al., 2024) with a LDM as $\hat{x}=\mathcal{G}(\hat{Y},E_{T5})$.
+In the inference phase, AudioLDM2 approximates LOA feature by the given condition as $\hat{Y}=\mathcal{M}(C)$ using a fine-tuned GPT-2 model [47].
+Then generates audios conditioned on the estimated LOA feature $\hat{Y}$ and an extra text embedding $E_{T5}$ from a FLAN-T5 [48] with a LDM as $\hat{x}=\mathcal{G}(\hat{Y},E_{T5})$.
 We denote the conditional embeddings in AudioLDM2 as $E=\{\hat{Y},E_{T5}\}$, therefore, the generative process becomes $\hat{x}=\mathcal{G}(E)$.
@@
 Therefore, we can derive the distribution of $z_{t}$ given $z_{0}$ as $q(z_{t}|z_{0})=\sqrt{\gamma_{t}}z_{0}+\sqrt{1-\gamma_{t}}\epsilon_{t}$, where $\gamma_{t}=\prod_{t=1}^{t}1-\beta_{t}$.
-The LDM learns a backward transition $\epsilon_{\theta}(z_{t},t)$ from the prior distribution $N(0,I)$ to the data distribution $z$, that predicts the added noise $\epsilon_{t}$ (Ho et al., 2020).
-Following the objective function of denoising diffusion probabilistic models (DDPM) (Ho et al., 2020), the objective function for training AudioLDM2 is
-$\text{min}_{\theta}\mathcal{L}_{DPM}=\text{argmin}_{\theta}[\mathbb{E}_{z_{0},E,t\sim\{1,...,T\}}||\epsilon_{\theta}(z_{t},E,t)-\epsilon_{t}||_{2}^{2}]$ (2)
-To reduce computational demands on inference, AudioLDM2 uses denoising diffusion explicit models (DDIM) (Song et al., 2020), which provides an alternative solution and enables significantly reduced sampling steps with high generation quality.
-The DDIM reverse diffusion process is
-$z_{t-1}=\sqrt{\gamma_{t-1}}(\frac{z_{t}-\sqrt{1-\gamma_{t}}\epsilon_{\theta}(z_{t},E,t))}{\sqrt{\gamma_{t}}})+\sqrt{1-\gamma_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(z_{t},E,t)+\sigma_{t}\epsilon_{t}$ (3)
-We can revise a deterministic mapping between $z_{0}$ and its latent state $z_{T}$ once the model is trained (Dhariwal & Nichol, 2021; Yang et al., 2023) by the following equation
+The LDM learns a backward transition $\epsilon_{\theta}(z_{t},t)$ from the prior distribution $N(0,I)$ to the data distribution $z$, that predicts the added noise $\epsilon_{t}$ [49].
+Following the objective function of denoising diffusion probabilistic models (DDPM) [49], the objective function for training AudioLDM2 is
+$\min\limits_{\theta}\mathcal{L}_{\text{DPM}}=\min_{\theta}[\mathbb{E}_{z_{0},E,t\sim\{1,..,T\}}||\epsilon_{\theta}(z_{t},E,t)-\epsilon_{t}||_{2}^{2}]$ (2)
+To reduce computational demands on inference, AudioLDM2 uses denoising diffusion explicit models (DDIM) [50], which provides an alternative solution and enables significantly reduced sampling steps with high generation quality.
+We can revise a deterministic mapping between $z_{0}$ and its latent STATE $z_{T}$ once the model is trained [51, 52] by the equation
 ${}\frac{z_{t+1}}{\sqrt{\gamma_{t+1}}}-\frac{z_{t}}{\sqrt{\gamma_{t}}}=(\sqrt{\frac{1-\gamma_{t+1}}{\gamma_{t+1}}}-\sqrt{\frac{1-\gamma_{t}}{\gamma_{t}}})\epsilon_{\theta}(z_{t},E,t)$ (3)
-4 Method
+Figure 1: Overview of SoundMorpher pipeline, where the snowflake represents the model parameters are frozen.
+IV Method
 Given a source and target audio pair $\{x^{(0)},x^{(1)}\}$, sound morphing aims to generate intermediate sounds $x^{(\alpha(t))}$ between the audio pair given morph factors $\alpha\in[0,1]$.
 To account for the variation of $\alpha$ over time in Equation 1, we discretize the function $\alpha(t)$ where $t\in[0,T]$ into $N$ elements, resulting in a morphed sequence of sounds $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ based on $\{\alpha_{i}\}^{N}_{i=1}$.
-According to the smoothness criteria proposed by Caetano & Osaka (2012), the desired sound morphing technique should have smooth linear perceptual stimuli when the morph factor $\alpha$ varies in the sequence $\{\alpha_{i}\}^{N}_{i=1}$.
-Therefore, we define $p_{i}$ to represent the perceptual stimuli of the morphed audio $x^{(\alpha_{i})}$ given morph factor $\alpha_{i}$.
+According to the smoothness criteria proposed by [16], a desired sound morphing results should exhibit a linear perceptual change as the morph factor $\alpha$ transitions through the sequence ${\alpha_{i}}^{N}_{i=1}$.
+Therefore, we define $p_{i}$ to represent the perception of the morphed audio $x^{(\alpha_{i})}$ given morph factor $\alpha_{i}$.
 However, the relationship $\mathcal{P(\cdot)}$ between morph factor $\alpha$ and perceptual stimuli $p$ is intractable.
-Our goal is to find a discrete morph factor sequence $\{\alpha_{i}\}^{N}_{i=1}$ such that for each transition, the perceptual stimuli difference $\Delta p$ is a constant value.
-Therefore, we formulate the problem as
-$p_{i+1}-p_{i}\equiv\mathcal{P}(x^{(\alpha_{i+1})})-\mathcal{P}(x^{(\alpha_{i})})=\Delta p,~{}~{}i\in[1,...,N-1]~{}$ (5)
-This formulation is a refined sound morphing problem where, rather than controlling morph factor $\alpha$, we control the constant perceptual stimuli difference $\Delta p$ to find the optimal trajectory with morph factors $\{\alpha_{i}\}^{N}_{i=1}$ that will achieve perceptually uniform sound morphing  [^1] See Appendix 7.1 for overall SoundMorpher pseudo algorithm pipeline and further implementation details..
-In Section 4.1 we introduce feature interpolation and model adaption with a pre-trained AudioLDM2.
-This method allows high-quality intermediate morph results to be obtained by controlling morph factor $\alpha$.
-To achieve perceptually uniform sound morphing as in Equation 5, we explore an explicit connection $\mathcal{P(\cdot)}$ between perceptual stimuli $p$ and morph factor $\alpha$ in Section 4.2.
-In Section 4.3, we provide extensions of our method on the different morphing methods discussed in Section 3.1 to show the advantages of perceptually uniform sound morphing.
-4.1 Feature interpolation and model adaption
+Our objective is to determine a discrete sequence of morph factors $\{\alpha_{i}\}^{N}_{i=1}$ such that the perceptual stimulus difference $\Delta p$ remains constant for each transition.
+We formulate the problem as
+$p_{i+1}-p_{i}\equiv\mathcal{P}(x^{(\alpha_{i+1})})-\mathcal{P}(x^{(\alpha_{i})})=\Delta p~{}$ (4)
+where $i\in[1,...,N-1]$.
+This formulation is a refined sound morphing problem where, rather than controlling morph factor $\alpha$, we control the constant perception difference $\Delta p$ to find the optimal trajectory with morph factors $\{\alpha_{i}\}^{N}_{i=1}$ that will achieve perceptually uniform sound morphing.
+The overall pipeline of SoundMorpher is presented in Fig.
+1 and Algorithm 1.
+Algorithm 1 Overall pipeline of SoundMorpher
+0:  A pre-trained AudioLDM2 pipeline including a pre-trained VAE with a encoder $g_{\theta}$ and a decoder $g_{\phi}$, a pre-trained latent diffusion model $\epsilon_{\theta}$, a pre-trained T5 model $f_{\phi}$, and a pre-trained GPT-2 model $f_{\varphi}$.
+Learning rates $\eta_{1},\eta_{2}$.
+Source and target audios $x^{(0)}$ and $x^{(1)}$.
+An initial text prompts $y$.
+perceptually uniform interpolation number $N$.
+tolerance error for binary search $\epsilon_{tol}$.
+Number of training steps for text inversion for conditional embedding optimization $T_{inv}$.
+Number of training steps for model adaptation $T_{adapt}$.
+LoRA rank $r$, step number of DDIM $T$.
+0:  Start morph factor $\alpha_{start}=0$, end morph factor $\alpha_{end}=1$. Start perceptual point $p_{start}=[0,1]$, and end perceptual point $p_{end}=[1,0]$.
+  Initialize: $z_{0}^{(0)}=g_{\theta}(x_{0}^{(0)})$, $z_{0}^{(1)}=g_{\theta}(x_{0}^{(1)})$; $E^{0}=[f_{\phi}(y),f_{\varphi}(y)]$, $E^{1}=[f_{\phi}(y),f_{\varphi}(y)]$;
+  # Step 1: Text-guided conditional embedding optimization
+  for $i$ from $1$ to $T_{inv}$ do
+     Randomly sample time step $t$ and random noise $\epsilon_{t}\sim N(0,I)$.
+     Adding noise to data $z_{t}^{(0)}\leftarrow\sqrt{\gamma_{t}}z_{0}^{(0)}+\sqrt{(1-\gamma_{t})}\epsilon_{t},$ $z_{t}^{(1)}\leftarrow\sqrt{\gamma_{t}}z_{0}^{(1)}+\sqrt{(1-\gamma_{t})}\epsilon_{t}$.
+     $E^{(0)}\leftarrow E^{(0)}-\eta_{1}\nabla_{E^{(0)}}\mathcal{L}_{\text{DPM}}(z_{0}^{(0)},E^{(0)};\theta)$.
+     $E^{(1)}\leftarrow E^{(1)}-\eta_{1}\nabla_{E^{(1)}}\mathcal{L}_{\text{DPM}}(z_{0}^{(1)},E^{(1)};\theta)$.
+  end for
+  # Step 2: Model adaptation with LoRA.
+  for $i$ from $1$ to $T_{adapt}$ do
+     Model adaptation with LoRA according to Equation 6 and Equation 7 with $\eta_{2}$ learning rate.
+  end for
+  # Step 3: Perceptual-uniform binary search with constant SPDP increment
+  Obtaining initial latent states $z_{T}^{(0)}$ and $z_{T}^{(1)}$ by Equation 3
+  $p_{list}\leftarrow$ ConstantSPDP$(N,p_{start},p_{end})$
+  $\alpha_{list}\leftarrow\text{BinarySeach}(\alpha_{start},\alpha_{end},p_{list},\epsilon_{tol})$
+  # Interpolate embeddings and latent state
+  for $\alpha$ in $\alpha_{list}$ do
+     $E^{(\alpha)}\leftarrow(1-\alpha)E^{(0)}+\alpha E^{(1)}$
+     $z_{T}^{(\alpha)}\leftarrow\frac{sin(1-\alpha)w}{sinw}z_{T}^{(0)}+\frac{sin\alpha w}{sinw}z_{T}^{(1)}$
+     for $t$ from T todo
+        $z_{t-1}^{(\alpha)}\leftarrow\sqrt{\gamma_{t-1}}(\frac{z_{t}-\sqrt{1-\gamma_{t}}\hat{\epsilon}_{\theta}^{(t)}(z_{t},E^{(\alpha)})}{\sqrt{\gamma_{t}}})+\sqrt{1-\gamma_{t-1}}\hat{\epsilon}_{\theta}^{(t)}(z_{t},E^{(\alpha)})$
+     end for
+  end for
+  $x^{(\alpha)}\leftarrow\text{vocoder}(g_{\phi}(z_{0}^{(\alpha)}))$
+  return  $\{x^{(\alpha)}\}_{\alpha\in\alpha_{list}}$
+In Section IV-A and Section IV-B, we introduce feature interpolation and model adaptation with a pre-trained AudioLDM2.
+This method allows high-quality intermediate morph results to be obtained by interpolating semantic conditional embeddings of source and target sounds according to morph factor $\alpha$.
+In Section IV-C, we explore an explicit connection $\mathcal{P(\cdot)}$ between perceptual stimuli $p$ and morph factor $\alpha$ to achieve perceptually uniform sound morphing, as in Equation 4.
+This approach enables us to refine the morph factors using binary search to achieve the desired perceptual differences $\Delta p$ for each transition.
+In Section IV-D, we provide extensions of our method on the different morphing methods discussed in Section III-A to show the advantages of perceptually uniform sound morphing.
+IV-A Feature Interpolation
+Similar to other sound morphing methods, SoundMorpher achieves sound morphing by interpolating features between source and target audios.
+In this section, we introduce a method to bridge the source and target audios based on semantic conditional embeddings in a pre-trained AudioLDM2, leading to a controllable sound morphing method by a morph factor $\alpha$.
 Interpolating optimized conditional embeddings.
-We first introduce text-guided conditional embedding optimization strategy under a pre-trained AudioLDM2, which retrieves corresponding conditional embeddings $E$ of the given audio data.
-As mentioned in Section 3.2, AudioLDM2 accepts two conditional inputs: LOA feature $Y$ and text embedding $E_{T5}$.
+We first introduce a text-guided conditional embedding optimization strategy under a pre-trained AudioLDM2 [42], which retrieves corresponding conditional embeddings $E$ of the given audio data.
+This optimization strategy enables our sound morphing method to handle diverse types of sounds without requiring the model to be retrained on specific datasets.
+As mentioned in Section III-B, AudioLDM2 accepts two conditional inputs: LOA feature $Y$ and text embedding $E_{T5}$.
 We denote $E=\{Y,E_{T5}\}$ as the overall conditional embedding inputs for AudioLDM2.
 The LOA feature $Y$ is an abstraction of audio data which is semantically structured, and $E_{T5}$ captures sentence-level of representations.
-To retrieve corresponding conditional embeddings of the given audio data, we first obtain the latent variables $z_{0}^{(0)}$ and $z_{0}^{(1)}$ of audio $x^{(0)}$ and $x^{(1)}$ from the pre-trained VAE in AudioLDM2 pipeline.
-We initialize a simple common text prompt (e.g., ‘An audio clip of sound’) as a text guidance condition $C$ to obtain $E$ by GPT-2 encoder and FLAN-T5 encoder in AudioLDM2 pipeline, respectively, as $E^{(0)}$ and $E^{(1)}$.
-Instead of optimizing the model parameters, we freeze the model parameters and optimize the conditional embedding $E^{(0)}$ and $E^{(1)}$ by the denoising objective function in Equation 2
-$E^{(0)}=\text{arg min}_{E}\mathcal{L}_{DPM}(z_{0}^{(0)},E;\theta)\text{ and }E^{(1)}=\text{arg min}_{E}\mathcal{L}_{DPM}(z_{0}^{(1)},E;\theta)$ (6)
+$\begin{split}E^{(0)}=\min_{E}\mathcal{L}_{\text{DPM}}(z_{0}^{(0)},E;\theta)\text{ and }\\ E^{(1)}=\min_{E}\mathcal{L}_{\text{DPM}}(z_{0}^{(1)},E;\theta)\end{split}$ (5)
 The optimized conditional embeddings $E^{(0)}$ and $E^{(1)}$ fully encapsulate the abstract details of audios $x^{(0)}$ and $x^{(1)}$.
@@
 Interpolating latent state.
-The conditional embedding represents the conceptual abstract of audio data.
-However, we also wish to smoothly morph the content of the audio pair.
-Following Song et al.
-(2020) and Yang et al.
-(2023), we smoothly interpolate between $z_{0}^{(0)}$ and $z_{0}^{(1)}$ by spherical linear interpolation (slerp) to their starting noise $z_{T}^{(0)}$ and $z_{T}^{(1)}$ and further obtained the interpolated latent state $z_{T}^{(\alpha)}:=\frac{\text{sin}(1-\alpha)\omega}{\text{sin}\omega}z_{T}^{(0)}+\frac{\text{sin}\alpha\omega}{\text{sin}\omega}z_{T}^{(1)}$, where $\omega=\text{arccos}(\frac{z_{T}^{(0)\intercal}z_{T}^{(1)}}{||z_{T}^{(0)}||||z_{T}^{(1)}||})$.
+The conditional embedding represents the conceptual abstract of audio.
+We also wish to smoothly morph the content of the audio pair.
+Following [50], we smoothly interpolate between $z_{0}^{(0)}$ and $z_{0}^{(1)}$ by spherical linear interpolation (slerp) to their starting noise $z_{T}^{(0)}$ and $z_{T}^{(1)}$ and further obtained the interpolated latent state $z_{T}^{(\alpha)}:=\frac{\text{sin}(1-\alpha)\omega}{\text{sin}\omega}z_{T}^{(0)}+\frac{\text{sin}\alpha\omega}{\text{sin}\omega}z_{T}^{(1)}$, where $\omega=\text{arccos}(\frac{z_{T}^{(0)\intercal}z_{T}^{(1)}}{||z_{T}^{(0)}||||z_{T}^{(1)}||})$.
 The denoised latent variable $z_{0}^{(\alpha)}$ is obtained by applying a diffusion denoising process on the interpolated starting noise $z_{T}^{(\alpha)}$ and conditioning on the interpolated conditional embedding $E^{(\alpha)}$.
 The final morphed audio result $x^{(\alpha)}$ is obtained from $z_{0}^{(\alpha)}$ by the VAE decoder and a vocoder.
-Model adaption.
-Model adaptation helps to limit the degree of morphed variation by suppressing high-density regions that not related to the given inputs (Yang et al., 2023).
-We use LoRA (Hu et al., 2021) to inject a small amount of trainable parameters for efficient model adaptation.
+IV-B Model adaptation
+Model adaptation helps to limit the degree of morphed variation by suppressing high-density regions that not related to the given inputs [52].
+To enhance the output quality of source and target audios.
+We follow [52] and use LoRA [53] to inject a small amount of trainable parameters for efficient model adaptation.
 We fine-tune AudioLDM2 with LoRA trainable parameters using $z^{(0)}$ and $z^{(1)}$.
-See Appendix 7.2 for details.
-4.2 perceptually uniform Sound Morphing
+Compared to vanilla fine-tuning approaches, LoRA has advantages in training efficiency with injecting small amount of trainable parameters.
+The model adaptation can be defined as
+$\begin{split}\min_{\Delta\theta^{\prime}}\mathcal{L}_{\text{DPM}}(z_{0}^{(0)},E^{(0)};\theta+\Delta\theta^{\prime})&+\\ \min_{\Delta\theta^{\prime}}\mathcal{L}_{\text{DPM}}(z_{0}^{(1)},E^{(1)};\theta+\Delta\theta^{\prime})&\text{ s.t.
+rank}(\Delta\theta^{\prime})=r\end{split}$ (6)
+where $r$ represents LoRA rank.
+To achieve high text alignment on inference, we use another LoRA parameters $\Delta\theta_{0}$ with rank $r_{0}$ to perform bias correction
+$\begin{split}\min_{\Delta\theta_{0}}\mathcal{L}_{\text{DPM}}(z_{0}^{(0)},\varnothing;\theta+\Delta\theta_{0})&+\\ \min_{\Delta\theta_{0}}\mathcal{L}_{\text{DPM}}(z_{0}^{(1)},\varnothing;\theta+\Delta\theta_{0})&\text{ s.t.
+rank}(\Delta\theta_{0})=r_{0}\end{split}$ (7)
+During inference, with $\theta^{\prime}=\theta+\Delta\theta^{\prime}$ and $\theta_{0}=\theta+\Delta\theta_{0}$, the noise prediction becomes
+$\hat{\epsilon}_{\theta}(z_{t},t,E):=w\epsilon_{\theta^{\prime}}(z_{t},t,E)+(1-w)\epsilon_{\theta_{0}}(z_{t},t,\varnothing).$ (8)
+Although [52] provide a heuristic suggestion for setting the LoRA rank value in the image morphing task; however, we further investigate the relationship between LoRA rank $r$ and method performance in our emprical experiment.
+IV-C Perceptually Uniform Sound Morphing
+Although interpolating semantic features between the source and target audios enables a smooth transition by leveraging the inherent structure of the semantic space, our ultimate goal is to explore the complex relationship between sound perception and the morph factor to further enhance the smoothness of the morphing sequence.
+In this section, we propose a method to further refine the morphed results by fixing the perception difference for each transition as constant as in Equation 4.
 Sound perceptual distance proportion (SPDP).
@@
 Our goal is to establish an objective quantitative metric that links $p_{i}$ and $x^{(\alpha_{i})}$ as in Equation 4.
-This metric should satisfy two key conditions: (1) the output $p$ should increase monotonically as $\alpha$ increases; (2) it should accurately represent perceptual differences between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$, ensuring a smooth transition through intermediate states.
-Therefore, we propose the sound perceptual distance proportion between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$.
+This metric should satisfy two key conditions: (1) the output $p$ should increase monotonically as $\alpha$ increases; (2) it should accurately represent sound perception and audio descriptive difference between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$, ensuring a smooth transition through intermediate states.
+To this end, we propose the sound perceptual distance proportion between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$.
 We define $p_{i}\in\mathbb{R}^{2}$ as a 2D vector to represent the perceptual proximity of $x^{(\alpha_{i})}$ to both $x^{(0)}$ and $x^{(1)}$.
-Instead of extracting numerous audio features through traditional signal processing techniques, we use Mel-scaled spectrogram to capture perceptual and semantic information on audio.
-Mel-spectrogram (Tzanetakis & Cook, 2002) provides a pseudo-3D representation of audio signals, with one axis representing time and the other representing frequency on the Mel scale (Stevens et al., 1937), while the values denote the magnitude of each frequency at specific time points.
-The advantage of using Mel-spectrogram lies in the Mel filter banks, which map frequencies to equal pitch distances that correspond to how humans perceive sound (Sturm, 2013; Müller, 2015).
-Denoting $x_{mel}^{(\alpha_{i})}$ as the Mel-spectrogram of audio $x^{(\alpha_{i})}$, the SPDP $p_{i}$ between two endpoint audios $x^{(0)}$ and $x^{(1)}$ given $\alpha_{i}$ is defined as
-$p_{i}=[\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}},\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}]$ (7)
+Instead of extracting numerous audio features through traditional signal processing techniques, we use log magnitude Mel-scaled spectrogram to capture perceptual information and descriptive aspects of audio, aligning well with the three objective criteria outlined in Section III-A.
+The log Mel-spectrogram [54] provides a pseudo-3D representation of audio signals, with one axis representing time and the other representing frequency on the Mel scale [55], while the values denote the magnitude of each frequency at specific time points.
+This visual representation enables humans to intuitively interpret audio characteristics and descriptive information through its structured depiction of temporal and spectral features.
+Another advantage of using Mel-spectrogram lies in the Mel filter banks, which map frequencies to equal pitch distances that correspond to how humans perceive sound [56, 57].
+A logarithmic scaling on the frequency further addresses the imbalance in human perception, which is inherently logarithmic and tends to disproportionately affect low and high-frequency regions.
+This approach aligns with the principles of psychoacoustics, which suggest that human auditory perception is better represented on a logarithmic frequency scale [58, 59].
+Denoting $x_{mel}^{(\alpha_{i})}$ as the log magnitude Mel-spectrogram of audio $x^{(\alpha_{i})}$, the SPDP point $p_{i}$ of $x^{(\alpha_{i})}$ between $x^{(0)}$ and $x^{(1)}$ given $\alpha_{i}$ is defined as
+$p_{i}=$ $[\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}},$ (9)
+$\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}]$
+Equation 9 calculates the distance proportion of how the morphed sample $x^{(\alpha_{i})}$ closes between source $x^{(0)}$ and target $x^{(1)}$, meanwhile the L2 norm measures the distance between two log Mel spectrograms.
+Therefore, the summation of the two elements in $p$ are always equals to 1.
+Furthermore, the first element of $p_{i}$ is strictly monotonically increasing with $\alpha$ while the second element is strictly monotonically decreasing, providing a clear and interpretable relationship between the parameter $\alpha$ and the perceptual proximity of $x^{(\alpha_{i})}$ to $x^{(0)}$ and $x^{(1)}$.
+In summary, the log Mel-spectrogram effectively captures both perceptual and descriptive audio information, offering an elegant solution to satisfy the criterions of smoothness, correspondence, and intermediateness proposed by [16].
 Binary search with constant SPDP increment.
-To produce a perceptually smooth morphing trajectory with a constant perceptual stimuli increment, we use binary search to seek the corresponding $\{\alpha_{i}\}^{N}_{i=1}$ based on a constant $\Delta p$.
-The target SPDP sequence $\{p_{i}\}_{i=1}^{N}$ is obtained by an interpolation $p_{i}=(1-\frac{i-1}{N-1})p^{(0)}+\frac{i-1}{N-1}p^{(1)}$, where the two endpoints are $p^{(0)}=[0,1]^{T}$ and $p^{(1)}=[1,0]^{T}$.
-See Algorithm 2 in Appendix 7.3 for detail pseudo algorithm.
-4.3 Controllable Sound Morphing with Discrete $\alpha$ Series
-By controlling the discrete morph factor sequence $\{\alpha_{i}\}^{N}_{i=1}$ to produce a morphed series $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ with constant $\Delta p$, we achieve three typical morphing methods as follows.
+To produce a perceptually smooth morphing trajectory with a constant perceptual stimuli increment, we use binary search to search for the corresponding $\{\alpha_{i}\}^{N}_{i=1}$ based on a constant $\Delta p$ as in Algorithm 2.
+The target SPDP sequence $\{p_{i}\}_{i=1}^{N}$ is obtained by interpolation $p_{i}=(1-\frac{i-1}{N-1})p^{(0)}+\frac{i-1}{N-1}p^{(1)}$, where the two endpoints are $p^{(0)}=[0,1]^{T}$ and $p^{(1)}=[1,0]^{T}$.
+Algorithm 2 Binary search with constant $\Delta p$
+0:  $\alpha_{start}$: start alpha value; $\alpha_{end}$: end alpha value; $N$: interpolate number; source audio $x^{(0)}$; target audio $x^{(1)}$;
+0:  $p_{list}=[]$, $p_{start}=[0,1]^{T}$, $p_{end}=[1,0]^{T}$; $\alpha_{list}=[]$, $\alpha_{start}=0$, $\alpha_{end}=1$;
+  # Step 1: Obtain target SPDP points with constant $\Delta p$.
+  Procedure ConstantSPDP$(N,p_{start},p_{end})$
+  for $i$ from $1$ to $N-1$ do
+     $t\leftarrow\frac{i}{N-1}$
+     $p_{i}\leftarrow(1-t)\times p_{start}+t\times p_{end}$
+     $p_{list}\leftarrow p_{list}\cup[p_{i}]$
+  end for
+  # Step 2: Perform binary search given target SPDP points.
+  Procedure BinarySearch($\alpha_{start},\alpha_{end},x^{(0)},x^{(1)},p_{list},\epsilon_{tol}$)
+  $\alpha_{list}\leftarrow[\alpha_{start}]$, $\alpha_{cur}\leftarrow\alpha_{start}$;
+  for $p_{i}$ from $p_{1}$ to $p_{N-2}$ do
+     $p_{target}\leftarrow p_{i}$
+     $\alpha_{t1}\leftarrow\alpha_{cur}$, $\alpha_{t2}\leftarrow\alpha_{end}$
+     $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
+     $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$
+     while $|p_{mid}-p_{i}|>\epsilon_{tol}$ do
+        if $p_{mid}>p_{target}$ then
+           $\alpha_{t2}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
+        else
+           $\alpha_{t1}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
+        end if
+        $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
+        $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$
+     end while
+     $\alpha_{list}\leftarrow\alpha_{list}\cup[\alpha_{mid}]$
+  end for
+  return  $\alpha_{list}$
+IV-D Controllable Sound Morphing with Discrete $\alpha$ Series
+Once a perceptually uniform discrete morph factor sequence $\{\alpha_{i}\}^{N}_{i=1}$ is obtained, we can further extend it to different sound morphing methods discussed in Section III-A.
 Static morphing.
@@
 We find the corresponding $\alpha$ value by the binary search with the target $p$ and further obtain a morphed result ${x^{(\alpha)}}$.
-Pseudocodes for static morphing are in Algorithm 3.
 Cyclostationary morphing.
 To produce $N$ perceptually uniform hybrid sounds between $x^{(0)}$ and $x^{(1)}$, we first obtain $N$ uniform interpolated SPDP points $\{p_{i}\}^{N}_{i=1}$.
-Then we find corresponding morph factors $\{\alpha_{i}\}_{i=1}^{N}$ and further obtain $N$ morphed results $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ as in Algorithm 4.
+Then we find corresponding morph factors $\{\alpha_{i}\}_{i=1}^{N}$ and further obtain $N$ morphed results $\{x^{(\alpha_{i})}\}_{i=1}^{N}$.
 Dynamic morphing.
-Dynamic morphing performs sound morphing over time, but one challenge is that if the morphing path fails to ensure perceptual intermediateness and content correspondence, the resulting sounds may exhibit perceptual discontinuities or unnatural intermediate stages.
-As in Algorithm 5, we obtain $N$ interpolated target SPDP points $\{p_{i}\}^{N}_{i=1}$ with $\Delta p$.
+Dynamic morphing performs sound morphing over time, one challenge is that if the morphing path fails to ensure perceptual intermediateness and audio description correspondence, the resulting sounds may exhibit perceptual discontinuities or unnatural intermediate stages.
+SoundMorpher produces perceptually uniform sound morphing sequences, which can avoid the problem when performing dynamic morphing.
+We obtain $N$ interpolated target SPDP points $\{p_{i}\}^{N}_{i=1}$ with $\Delta p$.
 The corresponding morph factors $\{\alpha_{i}\}_{i=1}^{N}$ are determined by binary search with the target SPDP points.
-Each morphed result $x^{(\alpha_{i})}$ contributes a segment of duration $\frac{T}{N}$, producing an audio segment $\tilde{x}^{(\alpha_{i})}$ according to index $i$.
-The final audio signal is obtained by concatenating these morphed segments, resulting in
-$[x_{0},x_{1},...,x_{T}]=\text{concat}(\tilde{x}^{(0)},\tilde{x}^{(\alpha_{1})},...,\tilde{x}^{(1)})$ (8)
-Table 1: Timbral morphing for musical instruments compared to baseline on different instruments.
+Each morphed latent variable $z^{(\alpha_{i})}$ contributes a segment of length $\frac{K}{N}$, where $K$ is the length of latent variables, producing a latent variable segment $\tilde{z}^{(\alpha_{i})}$ according to index $i$.
+The final morphed audio sample can be obtained by
+$x_{morph}=\text{vocoder}(\text{decoder}(\text{concat}(\tilde{z}^{(\alpha_{1})},...,\tilde{z}^{(\alpha_{N})})))$ (10)
+TABLE I: Timbral morphing for musical instruments compared to the baseline on different instruments.
 […]
@@
 Ours 3.29 19.73 0.233 0.023 $\pm$ 0.010 0.423 0.097
-5 Experiment
-In this section, we showcase three applications of SoundMorpher in real-world scenarios: Timbral morphing for musical instruments, Environmental sound morphing, and Music morphing.
-5.1 Evaluation Metric
+V Experiment
+In this section, we show three applications of SoundMorpher in real-world scenarios: Timbral morphing for musical instruments, Environmental sound morphing, and Music morphing.
+V-A Evaluation Metric
+We systematically verify SoundMorpher according to the criterions mentioned in Section III-A, with a series of adapted quantitative objective metrics as below
+  • Correspondence.
+We design a metric that computes absolute error for the mid-point MFCCs proportion, namely MFCCsE, for descriptive correspondence.
+Let $N$ be an odd integer, We define a series of perceptually uniform morphing results $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ with source and target audio $x^{(0)}$ and $x^{(1)}$, where $i$ in the range of $1$ to $N$.
+The MFCCsE is computed by
+    $(\displaystyle\text{MFCCs}_{\mathcal{E}}=\text{abs}(~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$11)
+    $\frac{||m^{(\frac{N+1}{2})}-m^{(0)}||_{2}}{||m^{(\frac{N+1}{2})}-m^{(0)}||_{2}+||m^{(\frac{N+1}{2})}-m^{(1)}||_{2}}$ 0.5-0.5)- 0.5 )
+    where $m^{(i)}$ represents the extracted MFCCs feature of the $i^{th}$ morphed results in the series $x^{(\alpha_{i})}$, $m^{(0)}$ and $m^{(1)}$ represents MFCCs feature of $x^{(0)}$ and $x^{(1)}$.
+This metric aims to evaluate spectrogram coherent of the midpoint result $x^{(\frac{N+1}{2})}$ between two end points $x^{(0)}$ and $x^{(1)}$, which MFCCs captures low-level of audio descriptive information [59, 60].
+Ideally, we wish the midpoint morphed result contains half-and-half audio description elements on two end points.
+The larger $\text{MFCCs}_{error}$ indicates the audio description consistency is far away than the midpoint (i.e., 0.5).
+We extract MFCCs feature with 13 coefficients to compute MFCCsE.
+    We use Fréchet audio distance (FAD) [61] and Fréchet inception distance (FID) [62] between morphed audios and sourced audios to verify semantic distribution similarity between morphed samples and sourced audios.
+Both metrics capture high-level of semantic audio descriptive information by extracting audio features from pretrained audio classification models.
+Given two sets of audio $X_{mrp}$ and $X_{src}$, where $X_{mrp}$ contains consecutive morphed samples as $X_{mrp}=\{x^{(\alpha_{i})}\}_{i=1}^{N}$, and $X_{src}$ contains sourced audio samples  [^1] In environmental sound morphing task, we set all sourced sounds in that class as $X_{src}$..
+We calculate FAD and FID values based on audio features extracted by pretrained audio classification models between $X_{src}$ and $X_{mrp}$  [^2] The FAD and FID metrics are obtained based on https://github.com/haoheliu/audioldm_eval.
 […]
-5.2 Timbral morphing for musical instruments
-Sound morphing can allow timbral morphing between the sound of two known musical instruments, creating sounds from unknown parts of the timbre space (McAdams, 2013; McAdams & Goodchild, 2017).
+  • Smoothness.
+We calculate the mean and standard deviation of CDPAM along with the morphing path to validate smoothness, as ${}_{mean}\pm\text{CDPAM}_{std}$, where $\text{CDPAM}_{mean}=\frac{1}{N-1}\sum^{N-1}_{i=1}\text{CDPAM}(x^{(\alpha_{i})},x^{(\alpha_{i+1})})$, and $\text{CDPAM}_{std}=\sqrt{\frac{1}{N-1}\sum^{N-1}_{i=1}(\text{CDPAM}(x^{(\alpha_{i})},x^{(\alpha_{i+1})})-\text{CDPAM}_{mean})^{2}}$.
+In timbre study, we define timbral distance as $\mathcal{L}_{2}^{timbre}$ by $\mathcal{L}_{2}^{timbre}=\frac{1}{N-1}\sum_{i=1}^{N-1}||q^{(\alpha_{i+1})}-q^{(\alpha_{i})}||_{2}$, where $q^{(\alpha_{i})}$ represents the corresponding timber point of $x^{(\alpha_{i})}$ in timbre space [66].
+  • Reconstruction of perceptual correspondence.
+Lastly, we denote CDPAME that calculate CDPAM between $\{x^{(0)},x^{(1)}\}$ and $\{\hat{x}^{(0)},\hat{x}^{(1)}\}$, where $\hat{x}$ represents resynthesized end points when $\alpha=0$ and $\alpha=1$.
+V-B Timbral Morphing for Musical Instruments
+Sound morphing can allow timbral morphing between the sound of two known musical instruments, creating sounds from unknown parts of the timbre space [67, 68].
 Timbral morphing for musical instruments involves transitioning between timbres of two different musical instruments to create a new sound.
@@
 Dataset.
-To ensure high quality of paired composition musical instrument on timbral study, we selected 22 paired musical instrument composition samples from demonstration pages of musical timbre transform projects, MusicMagus (Zhang et al., 2024) and Timbrer (Kemppinen P., 2020).
+To ensure high quality of paired compositions on timbral study, we selected 22 paired musical instrument composition samples from demonstration pages of musical timbre transform projects, MusicMagus [30] and Timbrer [69].
 The paired samples have durations varies from 5s to 10s, with 16.0kHz and 44.01kHz, which involve 5 groups of instrument pairs: 2 paired samples of piano-violin; 10 paired samples of piano-guitar; 1 paired sample of taiko-hihat; 1 paired sample of piano-organ, and 8 paired samples of harp-kalimaba.
 (a) SMT
-(b) SoundMopher (ours)
+(b) SoundMorpher (ours)
 Figure 2: Timbre space visualization of morph trajectories for piano-organ timbre morphing. Compared to SMT, SoundMorpher produces a smoother and continuous morph in the timbre space.
+Figure 3: Visualization of timbre morphing for musical instruments with $N=11$ by SoundMorpher compared to SMT.
 Baseline.
-We compare our method with Sound Morphing Toolbox (SMT) (Caetano, 2019), which is a set of Matlab functions targeting on musical instrument morphing that implement a sound morphing algorithm for isolated musical instrument sounds.
-Since SMT does not offer guidance for selecting perceptually uniform morph factors, we uniformly interpolate 11 morph factors in $[0,1]$.
+We compare with Sound Morphing Toolbox (SMT) [11], a set of Matlab functions targeting on musical instrument morphing.
+SMT implements a sound morphing algorithm for isolated musical instrument sounds.
+Since SMT performs sound morphing based on the linear assumption [16], we uniformly interpolate 11 morph factors in $[0,1]$.
 Results and analysis.
-We set $N=11$ for SoundMorpher with an initial prompt ‘a music composition by {instrument}’.
-The comparison of our method and the baseline on timbral morphing is in Table 1.
+The comparison of our method and the baseline on timbral morphing is in Table I.
 Overall, SoundMorpher demonstrates superior morphing quality compared to STM across various metrics, including audio quality, intermediateness, smoothness, and resynthesis quality   [^3] Since we perform timbral morphing within the same music composition, MFCCsE may not a suitable metric under the same musical content.
@@
 In contrast, SoundMorpher maintains robustness across different types of musical instruments, making it a more flexible and efficient solution for timbral morphing applications on different types of musical instruments.
+It is not surprising that SMT has poor performance in this experiment, since the traditional signal processing sound morphing methods are limited to complex or inharmonic sounds [12].
 Figure.
 2 provides a visualization of normalized timbre space, illustrating morphing trajectories generated by SMT and SoundMorpher.
-The timbre space is defined by three important timbral features: Log-Attack Time, Spectral Centroid, and Spectral Flux (McAdams et al., 1995; McAdams, 2013).
+The timbre space is defined by three important timbral features: Log-Attack Time, Spectral Centroid, and Spectral Flux [66, 67].
 The SMT trajectory shows distinct steps, indicating that the transitions between each intermediate sound are relatively abrupt.
@@
 This suggests that SoundMopher achieves a more continuous and natural-sounding morphing process, with each step being a smaller, more refined adjustment compared to SMT.
-Figure 7 and Figure 8 in the appendix provides additional visualization for this experiment.
-5.3 Environmental Sound Morphing
+Fig.
+3 provides spectrogram visualization of how SoundMorpher performs timbre morphing.
+V-C Environmental Sound Morphing
 Environmental sounds are used in video game production to provide a sense of presence within a scene.
@@
 Dataset.
-We use ESC50 (Piczak, 2015) which consists of 5-second recordings organized into 50 semantic classes which loosely arranged into 5 major categories.
+In this experiment, we firstly use ESC50 [70] to study the robuteness of SoundMorpher accross different categories of environmental sounds.
+Then we make a direct comparison with a concurrent method, MorphFader [13], based on environmental sounds provided in  [^4] https://pkamath2.github.io/audio-morphing-with-text/webpage/index.html.
+ESC50 [70] which consists of 5-second recordings organized into 50 semantic classes which loosely arranged into 5 major categories.
 We randomly select 4 major categories of scenarios to verify our method, including (1) Dog-Cat (animals voices), (2) Laughing-Crying baby (human sounds), (3) Church bells-Clock alarm (urban noise-interior sound), (4) Door knock-clapping (interior sounds-human sounds).
 Each category of scenarios contains 25 randomly selected audio pairs, thereby, 100 randomly paired samples in total.
-Table 2: Environmental sound morphing with different types of environmental sounds.
+To have a fair comparison, we also perform environmental sound morphing based on samples provided by MorphFader, which is sourced from AudioPairBank [71].
+Baseline.
+Due to lack of open-sourced direct comparison, we compare SoundMorpher with MorphFader [13] based on the criteria outlined in Section III-A.
+Although MorphFader didn’t release their codes, our evaluation system still can make a fair comparison based on their demonstrated examples.
+TABLE II: Environmental sound morphing with different types of environmental sounds.
 […]
@@
 Door knock $\leftrightarrow$ Clapping 21.36 0.083 10.85 76.35 1.594 0.428 $\pm$ 0.220 0.321
+Figure 4: Visualization of spectrogram for morphed results compred with MorphFader.
+SoundMorpher appears to provide a more seamless and stable morphing process, in which transitions are smoother and the spectral content is more consistent across the morphing stages.
+TABLE III: Comparison with MorphFader
+Method CDPAMT CDPAMmean±std MFCCsE
+MorphFader 0.972 0.243 $\pm$ 0.139 0.065
+SoundMorpher 0.935 0.226 $\pm$ 0.162 0.065
+Figure 5: Visualization of environmental sound morphing with $N=5$, from top to bottom: (1) church bells $\leftrightarrow$ clock alarm (2) crying baby $\leftrightarrow$ laughing (3) cat $\leftrightarrow$ dog (4) clapping $\leftrightarrow$ wood door knocking
 […]
-5.4 Music Morphing
+We then make a direct comparision with the concurrent method, MorphFader [13], based on their demonstrated morphed samples in Table III  [^5] Due to the demonstration page doesn’t provide original audio samples, we cannot compute FAD, FID and CDPAME for this comparison..
+SoundMorpher produces a smoother and intermediate morph than MorphFader with improvement on CDPAMT and CDPAMmean.
+Fig.
+4 provides a direct visual comparison of a morph sequence produced by MorphFader and our method.
+The yellow rectangles in Fig.
+4 highlight frequency band intensity cross morphing results.
+The intensity of the frequency bands within the yellow rectangle changes more abruptly for MorphFader, in contrast, SoundMorpher are more stable and consistent across time.
+This suggests that SoundMorpher maintains better spectral consistency during the morphing process, with smoother transitions between different timbral characteristics.
+As red rectangles indicate, MorphFader introduces more abrupt changes at the end of the morphing sequence.
+There is a noticeable shift in the pattern, indicating less smoothness on transition.
+In contrast, SoundMorpher shows a more gradual and consistent transition within the red rectangles.
+The spectral patterns remain more stable and exhibit smoother transitions towards the end of the morphing sequence.
+Overall, SoundMorpher appears to provide a more seamless and stable morphing process.
+The transitions are smoother, and the spectral content is more consistent across the morphing stages.
+V-D Music Morphing
 Film or game post-production often requires blending or fading between music tracks to seamlessly transition background music in between scenes.
@@
 Dataset.
-In this experiment, we randomly selected 50 sample pairs from 20 musical samples available on AudioLDM2 (Liu et al., 2024) demonstration page.
+Motivated by [30], we randomly selected 50 sample pairs from 20 high-quality musical samples available on AudioLDM2 [42] demonstration page.
 These 10-second music compositions that span different genres and feature both single or mixed musical instrument arrangements.
-Figure 2: Visualization of dynamic morphed music with $N=15$, source music and target music.
+Baseline.
+Due to lack of direct comparison on dynamic music morphing, we set the baseline as SoundMorpher without SPDP binary search, which achieves sound morphing by smoothly interpolate condition embeddings and latent states.
+We uniformly sample 15 morph factors in $\alpha\in[0,1]$ as the baseline for comparison.
+Figure 6: Visualization of dynamic morphed music for SoundMorpher and baseline with $N=15$, source music and target music.
 Results and analysis.
-In this experiment, we select an initial text prompt as ‘a sound clip of music composition’ to perform dynamic morphing.
-Even though this experiment contains morph complex music compositions with different music genres and music instruments, Table 3 shows our method still superiors on perceptual smoothly transiting source music to the target music and ensures correspondence, intermediateness and smoothness.
-Figure 2 provides strong visual evidence that the dynamic morphing method effectively transitions from the source to the target music while maintaining perceptual smoothness, correspondence, and intermediate transformations.
-The spectrogram illustrates that the morphed music transitions gradually, maintaining smooth spectral changes over time, which suggests the method successfully morphs the source into the target music.
-Table 3: Music morphing experimental results & ablation study for sound perceptual features, where N represents the number of components of PCA for reducing dimension of Mel-spectrogram.
+Even though this experiment contains morph complex music compositions with different music genres and music instruments, Table IV shows our method superiors on perceptual smoothly transiting source music to the target music and ensures correspondence, intermediateness and smoothness compared to the baseline, which further illustrates the advantage of the proposed algorithm of SPDP with binary serach.
+Fig.
+6 visually demonstrates the effectiveness of the proposed dynamic morphing method in transitioning smoothly from source to target music while ensuring perceptual consistency and intermediate transformations.
+Compared to the baseline results, which exhibit noticeable abrupt changes and spectral discontinuities, SoundMorpher produces gradual and smooth spectral transitions over time.
+This highlights that simply linearly interpolating semantic features, as done in the baseline method, fails to achieve truly smooth and perceptually consistent sound morphing.
+TABLE IV: Music morphing experimental results, baseline comparison & ablation study for sound perceptual features
 […]
+Baseline (Mel-Spec.) 0.099 10.82 58.20 0.938 0.067 $\pm$ 0.065 0.158
+SoundMorpher (Mel-Spec.) 0.047 9.85 56.25 0.847 0.068 $\pm$ 0.045 0.178
 Reduced Mel-Spec. (n=2) 0.187 10.31 58.57 0.793 0.056 $\pm$ 0.075 0.182
 Reduced Mel-Spec. (n=3) 0.151 10.76 59.39 0.779 0.055 $\pm$ 0.079 0.151
-Mel-Spec. 0.056 9.85 56.09 0.847 0.068 $\pm$ 0.045 0.178
 MFCCs 0.053 10.11 57.38 0.987 0.071 $\pm$ 0.050 0.156
-Spectral contras 0.066 10.54 58.44 0.863 0.061 $\pm$ 0.071 0.155
-Table 4: Mean opinion score study on environmental sound morphing and music morphing task.
-[…]
-Environmental sound morphing 3.78 $\pm$ 0.31 3.67 $\pm$ 0.40 3.57 $\pm$ 0.47 3.67 $\pm$ 0.39
-Music morphing 3.81 $\pm$ 0.39 3.55 $\pm$ 0.51 3.49 $\pm$ 0.48 3.62 $\pm$ 0.46
-5.5 Discussion
-Mean opinion score (MOS) study.
-We conducted a MOS study as a subjective evaluation for the correspondence, intermediateness, and smoothness of morphed results from SoundMorpher.
-The study involved 21 volunteers, and detailed methodology is in Appendix 9.
-As shown in Table 4, the results suggest that SoundMorpher is versatile, performing similarly well across both music and environmental sound morphing tasks, with no significant differences observed in the overall MOS.
-This consistency in scores indicates that SoundMorpher effectively handles the unique challenges posed by the distinct characteristics of music and environmental sounds, such as the continuous nature of music compared to the more segmented structure of environmental sounds.
-Despite the objective metric results showing clear differences between the two tasks, the human evaluation suggests that SoundMorpher remains robust across different sound types.
-One possible interpretation is that our objective metrics are more sensitive to variations in the measured aspects than participants.
-Table 5: Comparison with MorphFader
-Method CDPAMT CDPAMmean±std MFCCsE
-MorphFader 0.972 0.243 $\pm$ 0.139 0.065
-SoundMorpher 0.935 0.226 $\pm$ 0.162 0.065
-Model comparison.
-We compare SoundMorpher with a concurrent work, MorphFader (Kamath et al., 2024), based on the criteria outlined in Section 3.1.
-MorphFader relies on a pre-trained AudioLDM (Liu et al., 2023) and perform sound morphing by text instructions.
-We compare with 7 examples provided on the demonstration of MorphFader, that MorphFader uniformly samples $5$ morph factors in $[0,1]$, resulting in a morph path with $[0,0.25,0.5,0.75,1]$.
-In contrast, SoundMorpher finds $\alpha$ values according to constant $\Delta p$ with $5$ uniformly interpolate $p$ points by binary search.
-As in Table 5, SoundMorpher produces a smoother and perceptual intermediates morphing than MorphFader.
-See Appendix 10 for details.
+Spectral Contrast 0.066 10.54 58.44 0.863 0.061 $\pm$ 0.071 0.155
+$\star$ where $n$ represents the number of components of PCA for reducing dimension of Mel-spectrogram
+V-E Discussion And Ablation Study
 Ablation study on sound perceptual features.
 We verified the perceptual feature in SPDP in music morphing task.
-We select alternative music information retrieval features (MIR) including MFCCs with 13 coefficients (Logan et al., 2000), and spectral contrast (Jiang et al., 2002).
+We select alternative music information retrieval features (MIR) including MFCCs with 13 coefficients [60], and Spectral contrastt [72].
 We use principal component analysis (PCA) to reduce the dimensionality of Mel-spectrogram to further capture variation of spectral content over time, which is referred to as reduced Mel-Spec.
-(Stevens et al., 1937; Casey et al., 2008; Jiang et al., 2002).
-Table 3 shows performance comparisons of SoundMorpher with different features, SoundMorpher with Mel-spectrogram achieves better morphing quality in terms of correspondence and smoothness variation with smaller FAD, FD and CDPAMstd.
+[55, 73, 72].
+Table IV shows performance comparisons of SoundMorpher with different features, SoundMorpher with Mel-spectrogram achieves better morphing quality in terms of correspondence and smoothness variation with smaller FAD, FID and CDPAMstd.
 While Mel-spectrogram yields higher CDPAMmean, CDPAMT and MFCCsE compared to reduced Mel-Spec.
@@
 This suggests Mel-spectrogram, as a pseudo-3D representation, provide more perceptual and semantic information, which contributes to improve morph quality compared to higher-level features.
-Table 6: Experiment and ablation study results on music morphing.
+TABLE V: Experiment and ablation study results on music morphing.
 […]
@@
 Uninformative $\checkmark$ 0.057 10.37 55.89 1.036 0.074 $\pm$ 0.049 0.211
-Uninformative $\checkmark$ 0.056 9.85 56.09 0.847 0.068 $\pm$ 0.045 0.178
+Uninformative $\checkmark$ 0.047 9.85 56.25 0.847 0.068 $\pm$ 0.045 0.178
 Uninformative v.s.
@@
 And we use the given text prompts in AudioLDM2  [^6] https://audioldm.github.io/audioldm2/ as informative inital prompts.
-As in Table 6, informative initial text prompts may help with resynthesis quality and further improves morph correspondence.
+As in Table V, informative initial text prompts may help with resynthesis quality and further improves morph correspondence.
 Despite the improved resynthesis quality with informative initial text prompts, the results show a decline in morphing intermediateness and smoothness.
@@
 Inference steps.
-In our experiment, we follow the configuration of Zhang et al.
-(2024) and set DDIM steps to 100.
-To verify whether DDIM steps affect SoundMorpher performances, we compare with 20 DDIM steps in Table 6.
+In our experiment, we follow the configuration of [30] and set DDIM steps to 100.
+To verify whether DDIM steps affect SoundMorpher performances, we compare with 20 DDIM steps in Table V.
 Larger inference step seems to help for reconstruction quality and slightly imporves morph quality, however, performance differences between inference steps are not significant.
-This indicates SoundMorpher is robust for inference steps, and we extend this ablation study on environmental sound morphing task in Appendix 11.1.
-Thus, we suggest selecting a suitable DDIM step to trade-off overall binary search algorithm time-consuming and morph quality.
+This indicates SoundMorpher is robust for inference steps; therefore, we suggest selecting a suitable DDIM step to trade-off overall binary search algorithm time-consuming and morph quality.
+Figure 7: Failure cases for SoundMorpher with $N=5$.
+The source (left first) and target (right first) sounds that have significant semantic difference in contents, this leads SoundMorpher to produce abrupt transitions.
+TABLE VI: Ablation study on model adaptation with different LoRA rank $r$ on music morphing. Rank with -- represents results without LoRA model adaptation.
+[…]
+- 0.073 10.38 56.02 1.052 0.085 $\pm$ 0.054 0.198
+4 0.047 9.85 56.25 0.847 0.068 $\pm$ 0.045 0.178
+8 0.059 10.01 56.35 1.035 0.073 $\pm$ 0.051 0.180
+16 0.059 9.95 56.14 1.058 0.075 $\pm$ 0.052 0.169
+32 0.130 10.77 59.06 0.818 0.058 $\pm$ 0.082 0.158
+Ablation study on LoRA rank.
+In this experiment, we conduct an ablation study on model adaptation with LoRA on the task of music morphing.
+We test SoundMorpher with different LoRA ranks as well as SoundMorpher without model adaptation.
+Following [52], we train LoRA parameters for 150 steps with 1e-3 learning rate.
+We also set unconditional bias correction with $r_{0}=2$ for 15 steps with 1e-3 learning rate.
+Table VI shows the results of SoundMorpher with different rank size on model adaptation settings.
+According to Table VI, SoundMorher without model adaptation has obvious performance drop on morphing correspondence compares results with LoRA model adaptation.
+Even though a higher LoRA rank has a slight improvement on perceptual reconstruction quality, however, SoundMorpher with $r=32$ indicates poor correspondence with large MFCCsE and large smoothness variance CDPAMstd.
+This result indicates that SoundMorpher with higher LoRA rank not lead to a better morphing quality.
+When $r=4$, SoundMorpher achieves the best performance on smoothness, and correspondence compared to $r=8$, $r=16$ and $r=32$.
+Therefore, we suggest LoRA rank for model adaptation in SoundMorpher shouldn’t be too large.
 Limitations.
-The current implementation of SoundMorpher based on AudioLDM2 with 16.0kHz sampling rate, which may limit output audio quality.
+Although SoundMorpher is the first open-world sound morphing method, offering high-quality morphed results and broad applicability, we have identified the following limitations: The current implementation of SoundMorpher based on AudioLDM2 with 16.0kHz sampling rate, which may limit output audio quality.
 The conditional embeddings optimization only applies to sounds that can be produced by AudioLDM2.
-Sound examples that close to white noise, such as pure wind blowing used in MorphGAN (Gupta et al., 2023) are not easily generated by AudioLDM2, which makes the conditional embedding optimization produce low quality resynthesis sounds.
-We also observed that input sounds with a large semantic gap (e.g., Church bells-Clock alarm in Table 2) result in lower morphing quality.
-Furthermore, we observed when two audios exhibit significant temporal structure differences, such as environmental sounds, SoundMorpher may produce abrupt transitions, see Appendix 11.4 and Figure 6 for further details.
-6 Conclusion
-We propose SoundMorpher, a sound morphing method base on a pretrained diffusion model that produces perceptually uniform morphing trajectories.
-Unlike existing methods, we refined the sound morphing problem and explored an explicit connection between morph factor and perceptual stimuli of morphed results which offers better flexibility and higher morphing quality, making it adaptable to various morphing methods and real-world scenarios.
-We validate SoundMorpher by a series of objective quantitative metrics as well as mean opinion score study following criteria proposed by Caetano & Osaka (2012).
-These quantitative objective metrics may help to formalize future studies on sound morphing evaluation.
+Sound examples that close to white noise, such as pure wind blowing used in MorphGAN [12] are not easily generated by AudioLDM2, which makes the conditional embedding optimization produce low-quality resynthesis sounds.
+Although SoundMorpher produces high-quality sound morphing results, abrupt transitions can occur when the source and target sounds have significant temporal structure differences.
+A clear example of this is attempting to morph continuous environmental sounds with sounds that contain more silence, as in Fig.
+7.
+Environmental sounds often consist of discrete and temporally separated events, such as a dog barking or a cat meowing, which have distinct and abrupt characteristics.
+These are inherently different from the more continuous and harmonically structured nature of music, where elements blend more fluidly over time.
+As a result, creating smooth transitions between such disjointed environmental sounds can be more challenging, leading to the perception of more abrupt or less natural transitions in the morphing process.
+VI Conclusion
+We propose SoundMorpher, a open-world sound morphing method base on a pretrained diffusion model that produces perceptually uniform morph trajectories.
+Unlike existing methods, we refined the sound morphing problem and explored a more complex connection between morph factor and perception of morphed results which offers better flexibility and higher morphing quality, making it adaptable to various morphing methods and real-world scenarios.
+We validate SoundMorpher with a series of adapted objective quantitative metrics following three format sound morphing criterions proposed by [16].
+These quantitative metrics may help to formalize future studies for direct comparison on sound morphing evaluation.
 Furthermore, we demonstrated that SoundMorpher can be applied to wide range of real-world applications in our experiments and conducted in-depth discussions.
 SoundMorpher also has the potential to achieve voice morphing, as its foundational model AudioLDM2 supports speech generation; however, we leave this exploration for future work.
-Appendix
-7 SoundMorper implementation details
-This section provides further details on SoundMorpher pipeline and implementations.
-7.1 Overall pipeline of SoundMorpher
-The overall pseudo pipeline for SoundMorpher is provided in Algorithm 1, where the overall pipeline of SoundMorpher contains three parts: (1) Conditional embedding optimization.
-(2) Model adaptation.
-(3) perceptually uniform binary search with constant SPDP increment.
-Algorithm 1 Pipeline of SoundMorpher
-A pre-trained AudioLDM2 pipeline including a pre-trained VAE with a encoder $g_{\theta}$ and a decoder $g_{\phi}$, a pre-trained latent diffusion model $\epsilon_{\theta}$, a pre-trained T5 model $f_{\phi}$, and a pre-trained GPT-2 model $f_{\varphi}$.
-Learning rates $\eta_{1},\eta_{2}$.
-Source and target audios $x^{(0)}$ and $x^{(1)}$.
-An initial text prompts $y$.
-perceptually uniform interpolation number $N$.
-Tolerance error for binary search $\epsilon_{tolerance}$.
-Number of training steps for text inversion for conditional embedding optimization $T_{inv}$.
-Number of training steps for model adaptation $T_{adapt}$.
-LoRA rank $r$, Number of steps for DDIM $T$.
-Start morph factor $\alpha_{start}=0$, end morph factor $\alpha_{end}=1$. Start perceptual point $p_{start}=[0,1]$, and end perceptual point $p_{end}=[1,0]$.
-Initialize: $z_{0}^{(0)}=g_{\theta}(x_{0}^{(0)})$, $z_{0}^{(1)}=g_{\theta}(x_{0}^{(1)})$; $E^{0}=[f_{\phi}(y),f_{\varphi}(y)]$, $E^{1}=[f_{\phi}(y),f_{\varphi}(y)]$;
-# Step 1: Text-guided conditional embedding optimization.
-for $i$ from $1$ to $T_{inv}$ do
-     Randomly sample time step $t$ and random noise $\epsilon_{t}\sim N(0,I)$.
-     Adding noise to data $z_{t}^{(0)}\leftarrow\sqrt{\gamma_{t}}z_{0}^{(0)}+\sqrt{(1-\gamma_{t})}\epsilon_{t},z_{t}^{(1)}\leftarrow\sqrt{\gamma_{t}}z_{0}^{(1)}+\sqrt{(1-\gamma_{t})}\epsilon_{t}$.
-     $E^{(0)}\leftarrow E^{(0)}-\eta_{1}\nabla_{E^{(0)}}\mathcal{L}_{DPM}(z_{0}^{(0)},E^{(0)};\theta)$.
-     $E^{(1)}\leftarrow E^{(1)}-\eta_{1}\nabla_{E^{(1)}}\mathcal{L}_{DPM}(z_{0}^{(1)},E^{(1)};\theta)$.
-end for
-# Step 2: Model adaptation with LoRA.
-for $i$ from $1$ to $T_{adapt}$ do
-     Model adaptation with LoRA according to Equation 9 and Equation 10 with $\eta_{2}$ learning rate.
-end for
-# Step 3: Perceptual-uniform binary search with constant SPDP increment.
-Obtaining initial latent states $z_{T}^{(0)}$ and $z_{T}^{(1)}$ by Equation 4
-$p_{list}\leftarrow$ ConstantSPDP$(N,p_{start},p_{end})$triangleright Obtain target SPDP points by Algorithm 2
-$\alpha_{list}\leftarrow\text{BinarySeach}(\alpha_{start},\alpha_{end},p_{list},\epsilon_{tolerance})$
-for $\alpha$ in $\alpha_{list}$ do
-     $E^{(\alpha)}\leftarrow(1-\alpha)E^{(0)}+\alpha E^{(1)}$
-     $z_{T}^{(\alpha)}\leftarrow\frac{sin(1-\alpha)w}{sinw}z_{T}^{(0)}+\frac{sin\alpha w}{sinw}z_{T}^{(1)}$
-     for $t$ from T todo
-         $z_{t-1}^{(\alpha)}\leftarrow\sqrt{\gamma_{t-1}}(\frac{z_{t}-\sqrt{1-\gamma_{t}}\hat{\epsilon}_{\theta}^{(t)}(z_{t},E^{(\alpha)})}{\sqrt{\gamma_{t}}})+\sqrt{1-\gamma_{t-1}}\hat{\epsilon}_{\theta}^{(t)}(z_{t},E^{(\alpha)})$
-     end for
-end for
-$x^{(\alpha)}\leftarrow\text{vocoder}(g_{\phi}(z_{0}^{(\alpha)}))$triangleright Decode latent variable and obtain audio waveform by a vocoder.
-return $\{x^{(\alpha)}\}_{\alpha\in\alpha_{list}}$
-7.2 Model adaptation with LoRA
-In the task of image morphing, Yang et al.
-(2023) indicate adapting the model to the input pair helps to limit the degree of morphed variation by suppressing high-density regions that are not related to the given images.
-Compared to vanilla fine-tuning approaches, LoRA has advantages in training efficiency with injecting limited trainable parameters.
-The model adaptation can be defined as
-$\text{min}_{\Delta\theta^{\prime}}\mathcal{L}_{DPM}(z_{0}^{(0)},E^{(0)};\theta+\Delta\theta^{\prime})+\text{min}_{\Delta\theta^{\prime}}\mathcal{L}_{DPM}(z_{0}^{(1)},E^{(1)};\theta+\Delta\theta^{\prime})\text{ s.t.
-rank}(\Delta\theta^{\prime})=r$ (9)
-where $r$ represents LoRA rank.
-Unconditional Bias Correction. To achieve high text alignment during inference time, we use additional LoRA parameters $\Delta\theta_{0}$ with a small rank $r_{0}$ to perform bias correction, as
-$\text{min}_{\Delta\theta_{0}}\mathcal{L}_{DPM}(z_{0}^{(0)},\varnothing;\theta+\Delta\theta_{0})+\text{min}_{\Delta\theta_{0}}\mathcal{L}_{DPM}(z_{0}^{(1)},\varnothing;\theta+\Delta\theta_{0})\text{ s.t.
-rank}(\Delta\theta_{0})=r_{0}$ (10)
-During inference, with $\theta^{\prime}=\theta+\Delta\theta^{\prime}$ and $\theta_{0}=\theta+\Delta\theta_{0}$, the noise prediction becomes
-$\hat{\epsilon}_{\theta}(z_{t},t,E):=w\epsilon_{\theta^{\prime}}(z_{t},t,E)+(1-w)\epsilon_{\theta_{0}}(z_{t},t,\varnothing).$ (11)
-In our experiment, we set $r=4$ and $r_{0}=2$.
-Although Yang et al.
-(2023) provide a heuristic suggestion for setting the LoRA rank value for image morphing task, however, we further investigate the relationship between LoRA rank $r$ and method performance in Table 8 in sound morphing task and discussion in Appendix 11.2.
-7.3 perceptually uniform binary search with constant SPDP increment
-Algorithm 2 provides detail pesudocodes for how to implement perceptually uniform binary search with constant SPDP increment.
-This pseudo algorithm includes two steps, firstly, compute constant SPDP increment according to interpolte point number $N$ and obtain $N$ target SPDP points as $\{p_{i}\}_{i=1}^{N}$.
-Secondly, perform binary search to find correponding morph factor $\alpha$ series $\{\alpha_{i}\}_{i=1}^{N}$ according to $\{p_{i}\}_{i=1}^{N}$.
-Algorithm 2 Pseudo algorithm for perceptually uniform binary search with constant SPDP increment
-$\alpha_{start}$: starting alpha value; $\alpha_{end}$: ending alpha value; $N$: number of interpolations; source audio $x^{(0)}$; target audio $x^{(1)}$;
-$p_{list}=[]$; $p_{start}=[0,1]^{T}$; $p_{end}=[1,0]^{T}$; $\alpha_{list}=[]$;
-# Step 1: Obtain target SPDP points with constant increment.
-Procedure ConstantSPDP$(N,p_{start},p_{end})$
-for $i$ from $1$ to $N-1$ do $\triangleright$ Find target SPDP points with constant $\Delta p$.
-     $t\leftarrow\frac{i}{N-1}$;
-     $p_{i}\leftarrow(1-t)\times p_{start}+t\times p_{end}$
-     $p_{list}\leftarrow p_{list}\cup[p_{i}]$;
-end for
-# Step 2: Perform binary search given target SPDP points with constant $\Delta p$.
-Procedure BinarySearch($\alpha_{start},\alpha_{end},x^{(0)},x^{(1)},p_{list},\epsilon_{tolerance}$);
-$\alpha_{list}\leftarrow[\alpha_{start}]$;
-$\alpha_{cur}\leftarrow\alpha_{start}$;
-for $p_{i}$ from $p_{1}$ to $p_{N-2}$ do
-     $p_{target}\leftarrow p_{i}$
-     $\alpha_{t1}\leftarrow\alpha_{cur}$;
-     $\alpha_{t2}\leftarrow\alpha_{end}$;
-     $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$;
-     $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$triangleright Compute SPDP by Equation 7.
-     while $|p_{mid}-p_{i}|>\epsilon_{tolerance}$ do $\triangleright$ Perform binary search
-         if $p_{mid}>p_{target}$ then
-              $\alpha_{t2}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
-         else
-              $\alpha_{t1}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
-         end if
-         $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
-         $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$
-     end while
-     $\alpha_{list}\leftarrow\alpha_{list}\cup[\alpha_{mid}]$triangleright Append the reult to $\alpha_{list}$.
-end forreturn $\alpha_{list}$;
-Algorithm 3 Pseudo algorithm for static morphing
-Source audio $x^{(0)}$, target audio $x^{(1)}$, specific SPDP point $p$, tolerance error for binary search $\epsilon_{tol}$, number of steps for DDIM $T$, VAE decoder $g_{\phi}$.
-$\alpha_{start}=0,\alpha_{end}=1$<